### v1

# Are Users More Willing to Use Formally Verified Password Managers?

Abstract.
Formal verification has recently been increasingly used to prove the correctness and security of many applications. It is attractive because it can prove the absence of errors with the same certainty as mathematicians proving theorems. However, while most security experts recognize the value of formal verification, the views of non-technical users on this topic are unknown. To address this issue, we designed and implemented two experiments to understand how formal verification impacts users. Our approach started with a formative study involving 15 participants, followed by the main quantitative study with 200 individuals. We focus on the application domain of password managers since it has been documented that the lack of trust in password managers might lead to lower adoption. Moreover, recent efforts have focused on formally verifying (parts of) password managers. We conclude that formal verification is seen as desirable by users and identify three actional recommendations to improve formal verification communication efforts.
Password Manager, Usable Security, Formal Methods, Formal Verification, User Study, HCI
journal: JACMjournalvolume: 37journalnumber: 4article: 111publicationmonth: 8ccs: Security and privacy Usability in security and privacyccs: Security and privacy Formal methods and theory of security
1. Introduction
Formal verification has developed substantially in recent years and has been applied to many domains. Recent work in this area includes the formal verification of security properties in trusted execution environments (jangid2021towards, ), critical parts of the Linux kernel (nelson2020specification, ), the open-source TLS implementation used in numerous Amazon services (chudnov2018continuous, ), and even compilers (leroy2006formal, ; kumar2014cakeml, ). Formal verification works by rigorously proving the correctness of code against a formal specification and thus ensuring that software components behave as intended, even under adversarial conditions. An extensive survey conducted in 2020 (2020_expert_surveyFM, ) revealed that experts in formal methods widely believe these approaches offer enhanced code quality, strengthened cybersecurity, streamlined certification processes, and more manageable maintenance efforts. Although security experts generally recognize the value of formal verification, the views (or even awareness) of end-users are unknown (carreira2021exploring, ; carreira2022studying, ).
Previous research has explored user expectations and perceptions of security in areas such as differential privacy (cummings2021need, ) and cryptocurrency systems (mai2020user, ). Similarly, prior work on formal methods has also successfully evidenced trends, limitations, and future paths for formal methods (2020_expert_surveyFM, ). However, to the best of our knowledge, the impact of formal verification on end-user willingness to use formally verified products is still unexplored. Understanding users’ views on formal verification can help foster the adoption of formally verified code (with all the advantages verified code can provide) and aid developers in communicating with end-users and industry partners.
Our work addresses this gap in the literature by studying formal verification’s impact on users of a concrete application domain: Password Managers. Password Managers are applications that help users generate and manage their passwords. Most Password Managers offer password security features like password generation, multi-factor authentication, and secure storage with a primary password. [^1] Some related work (pearman2019people, ; ray2020older, ; chiasson2006usability, ) uses “master password” instead of “primary password”. However, recent work uses “primary password” (e.g., Firefox’s Password Manager (mozilla, )).
We chose this domain because text passwords are one of the most used security mechanisms, and Password Managers are an essential security tool to manage them (persistenceOfPasswords, ). However, previous work suggests that users have trust issues with Password Managers (ion2015no, ; ray2020older, ) and are reluctant to use them due to a lack of understanding of their security properties (pearman2019people, ). In parallel, recent work on formal verification has been applied to Password Managers (grilo2021towards, ; grilo22pwdgen, ). We designed and deployed two studies to understand users’ views on formal verification. The first is a formative interview study, asking users to use a Password Manager with verified components. The main study is a more extensive quantitative study that builds upon the formative and aims to explore some of the findings and themes identified in more detail.
Contributions.
Overall, our main contributions are:
Addressing a critical gap in the literature: Our research provides an empirical evaluation of the influence of formal verification on end-users within a specific application domain —- Password Managers. This work makes a contribution by addressing a gap in the formal methods literature – the study of end-users of formal verification.
Impact of formal verification on user adoption: Our findings suggest that formal verification has a positive effect on users’ willingness to adopt Password Managers. In both our formative user study and main study, participants showed an increased willingness to use Password Managers with formally verified components and identified formal verification as a desirable feature.
Identification of features for formal verification: We systematically identify a hierarchy of Password Manager features that users consider most critical for formal verification. We identify several critical features that are particularly relevant for participants. We argue that developers and practitioners should focus their formal verification efforts on features that users deem most critical for their trust and usage intentions, such as vault security and password generation.
Recommendations to improve formal verification communication efforts: Finally, we contribute three actionable recommendations for industry practitioners. We argue that practitioners can increase the appeal and effectiveness of verified security software by: i) addressing user-identified priorities, ii) enhancing transparency around the function and benefits of formal verification, and iii) engaging in broader educational initiatives to educate users about formal verification.
2. Scope and Research Questions
Our formative study motivates three questions. This study aims to test users’ perceptions of formal verification. To this end, we choose the domain of Password Managers and specifically address the following Research Questions:
RQ1.:
How does formal verification impact users’ willingness to use Password Managers?
RQ2.:
What features would users like to see formally verified in a Password Manager?
RQ3.:
Do users value the guarantees formal verification can provide in Password Managers?
Summary of Methods.
Figure 1. Overview of the Methodology
Our overarching goal is to understand how formal verification in Password Managers influences the decisions and trust of non-expert users.
The formative study, characterized by its qualitative and exploratory nature, focuses solely on RQ1. It aims to capture users’ initial reactions and understandings of formal verification in Password Managers through in-depth interviews and interactions with a Password Manager prototype. This prototype, an extended version of Bitwarden (bitwarden, ), incorporates a formal verification icon and explanations to educate users about formal verification’s role in enhancing Password Manager security. It gathers insights on general themes around Password Managers and formal verification, setting the groundwork for the subsequent larger-scale survey. The emphasis here is on identifying and exploring users’ perceptions, which directly contributes to addressing RQ1 by revealing how users view and understand formal verification in the context of Password Managers.
The main study builds on the findings from the formative study, expanding the scope to address all research questions mentioned in the introduction. With 200 participants, this larger-scale survey aims to quantitatively assess the broader implications of formal verification on users’ willingness to use Password Managers. This study aims to validate the themes identified in the formative study at a larger scale and further explore how formal verification impacts users’ choices and trust in Password Managers. We use a combination of statistical analysis and scenario-based questions to understand better the role of formal verification across different aspects of Password Manager usage and answer our research questions.
Crucially, we do not wish to ”explain” formal verification to our participants; we merely want to communicate its consequences and impact on the Password Manager. An overview of the methodology adopted can be seen in Figure 1.
Ethical Considerations.
In both studies, we did not collect any personal data. All participants were over 18 years old and were shown a consent form, the terms of which they had to accept before starting the survey. In the first author’s institution (the lead institution for this project), studies with these properties do not need to be submitted to the Ethics Committee (as confirmed by the Chair of the Ethics Committee). The local Ethics Committee recommends reviewing the information and checklists shown in the European Commission’s guide on ethical self-assessment. [^2] European Commission’s guide on ethical self-assessment can be accessed here https://ec.europa.eu/info/funding-tenders/opportunities/docs/2021-2027/common/guidance/how-to-complete-your-ethics-self-assessment_en.pdf. All studies described in this paper followed these guidelines. Co-authors from other institutions did not have access to individual survey data and only worked with aggregate statistics, as reported in this paper.
3. Background and Motivation
Formal verification.
In general, it is hard to build secure computer-based systems. Formal methods offer the promise of software that does not have exploitable bugs (fisher2017hacms, ) and has been used in a wide range of domains to prove the correctness and security of applications (jangid2021towards, ; nelson2020specification, ; grilo22pwdgen, ; chudnov2018continuous, ; ferreira2017certified, ; grilo2022verified, ). The idea of formal verification is mathematically modeling a system and then using formal methods to prove that the model satisfies specific properties or specifications. For example, Chudnov et al. (chudnov2018continuous, ) describe the development and operation of a continuously checked proof ensuring key properties of the TLS implementation used by many AWS services. Nelson et al. (nelson2020specification, ) describe their experience applying formal methods to a critical component in the Linux kernel, the just-in-time compilers. Their results show that building a verified component within a sizeable unverified system is possible with careful specification and proof strategy design. Another line of work investigates the application of formal methods to the specification and evaluation of password composition policies. Johnson et al. (johnson2020skeptic, ) introduce Skeptic, a toolchain that applies formal verification and power-law modeling to automatically select and justify password composition policies. Formally verified compilers are also available. An example is the CompCert (leroy2009formal, ) compiler, which compiles code from a large subset of the C programming language to PowerPC assembly code and guarantees that safety properties proved on the source code hold for the executable compiled code as well (leroy2009formal, ).
Password managers.
Passwords have become increasingly complex as most websites adopt strict password security policies.Previous work suggests users struggle with generating unique passwords (inglesant2010true, ), and password reuse is a problem—Pearman et al. (pearman2017let, ) show that 40% of users reuse 81–90% of their passwords. The basic features that all Password Managers provide to users are: (a) A secure encrypted vault to save passwords; (b) A password generator to generate unpredictable passwords that comply with complex password security policies. Moreover, most modern Password Managers also provide other features, including: (c) “Autofill” to automatically fill passwords in a website’s login screen; (d) Cloud synchronization to keep passwords safe and synchronized across devices. Despite being recommended by experts  (enisapassword, ), Password Manager usage is not widespread. Users complain about not knowing that Password Managers exist and not understanding how they work (pearman2019people, ; ion2015no, ; ray2020older, ; chiasson2006usability, ). Users also do not trust Password Managers (pearman2019people, ). As an intermediary tool, Password Managers must be reliable, consistent, and predictable. Previous work (chiasson2006usability, ; pearman2019people, ) recommends education as a tool for enhancing adoption and trust in Password Managers.
Usable security.
An effective security mechanism is one that is used correctly. Usable security can be traced back to Saltzer and Schroeder’s 1974 paper (saltzer1975protection, ) that introduces the term “psychological acceptability” for access-control systems. In the last decade, there has been a lot of work on usable security across many domains (mai2020user, ; pearman2019people, ; ion2015no, ; carreira2021towards, ). Prior work explored various dimensions of security warnings, from SSL warnings in web browsers to end-to-end encryption guarantees. A notable contribution is the series of studies by Felt et al. (felt2015improving, ) examining the impact of SSL warnings on user behavior within Google Chrome. Their work focused on experimenting with SSL warnings and assessing user adherence and comprehension. Similarly, Akhawe and Felt’s (akhawe2013alice, ) study provided evidence on the effectiveness of browser security warnings, revealing insights into how users perceive and react to these alerts in real-world settings. Other relevant, usable security work includes previous efforts to communicate about security using labels similar to “nutrition labels (kelley2009nutrition, ).” In this realm, previous research designed privacy labels for several domains, such as app stores (li2022understanding, ) and IoT devices (emami2020ask, ). Related work has also evaluated the effectiveness of commonly deployed password strength meters and found significant inconsistencies between meter classifications and actual password resistance to cracking: many passwords labeled as “strong” were, in fact, easily guessed, and some “weak” passwords resisted cracking (pereira2020evaluating, ).
Limitations of Existing Approaches.
A 2020 survey (2020_expert_surveyFM, ) of formal methods experts studied their views of formal methods. However, to the best of our knowledge, no previous work has focused on (non-expert) users of formally verified Password Managers, nor, more generally, on the impact that formal verification may have on end users. Previous work on formal verification of Password Managers is scarce, but efforts have been made to formally verify the random password generation algorithm of a Password Manager (grilo2021towards, ; grilo22pwdgen, ). However, these studies concentrated on the technical aspects of formal verification without examining its usability implications.
4. Formative Study
Our goal with this formative study was to understand users’ overall views on formal verification. To achieve this, we implemented two significant extensions to a Password Manager: a formal verification icon and an explanation.
4.1. Method
We created a proof-of-concept prototype by extending an established Password Manager’s – Bitwarden’s (bitwarden, ) browser extension – interface and testing it for user acceptance.
Formal verification icon.
Bitwarden uses icons to communicate with users. Similarly, we use an icon to communicate formal verification as icons are a good alternative to text to communicate with users (Wiedenbeck1999, ). The formal verification icon we chose is shown in Figure 2 and was placed on all features that could be formally verified: (a) Password vault, by the password field; (b) Primary password input box; (c) Password generator; (d) Clipboard, by its settings.
We chose green due to its association with safety in Anglo-American contexts (dix2004human, ) and derived the green used in the UI from Bitwarden’s blue (#175DDC ) and grey (#7C7C7C ), ending up with the shade (#0BDB0B ).
We designed the formal verification icon with Font Awesome 4333Font Awesome is fully open source and GPL font friendly. https://fontawesome.com, the same font used to design the existing Bitwarden’s iconography. We chose green due to its association with safety in Anglo-American contexts (dix2004human, ) and derived it from Bitwarden’s blue (#175DDC) and grey (#7C7C7C), ending up with the shade (#0BDB0B) using Adobe Color [^4] A palette tool referenced by Tidwell et al. (tidwell2010designing, ). https://color.adobe.com/create. The design process followed three phases: (1) Initial brainstorming with security researchers that resulted in 15 icon designs, later narrowed to 7. (2) Contextual tests of the icons within the PMs interface led to feedback on aesthetics, size, and positioning. (3) An external survey with 20 users with: (a) an initial preference test without explanation, (b) a preference test informed by the icon’s purpose (harley_2016, ), (c) final ratings in the icon’s intended context. A detailed description of the icon design process can be found in Appendix A.
Figure 2. Formal Verification icon
Descriptions of formal verification.
To explain the role of formal verification in the Password Manager, we initially identified all features amenable to formal verification, as well as the potential locations within the Password Manager interface for displaying the verification icon.
We developed the explanations in three phases: (1) We began with one-on-one discussions with formal verification researchers, gathering explanations, and finally, removing jargon. (2) The second iteration consisted of meeting with the team, gathering feedback , and applying it. (3) Finally, after another round of feedback, we reached a consensus and finalized the explanations. We compared the base Password Manager (i.e., without any extensions or formal verification) and the extended Password Manager (i.e., with interface updates). We performed two pilot interviews to refine the protocol and interview script.
Structure.
The user tests were divided into four parts. First, we provided users with a brief introduction to the study, and after we asked users to fill out the “Pre-Task Questionnaire” about their experience with Password Managers and demographics. All participants were asked to perform everyday tasks in a Password Manager (e.g., save a password). And finally, we ask them to fill out the “Final Questionnaire” . Each questionnaire was used as a base for semi-structured interviews. We asked follow-up questions about the participants’ answers. Data was collected in two ways: through questionnaires and observation. The extended Password Manager did not have formally verified features. As such, we included the following disclaimer: “This is a product in development, and some formally verified features are not fully implemented.”.
Recruitment and Participant demographics.
Our sample comprised 15 participants – 10 for the extended interface and 5 for the control condition. Participants were recruited through the authors’ network and received no compensation. Most of the participants (60%) had higher education. Of the 15 participants, only 2 had a technical background related to IT. The most frequent age group was 25-34 (40%). Overall, 60% of users were younger than 34, and 40% were at least 35 years old. Gender in our sample was evenly divided. Each interview took around 50 minutes, and 90% of participants were unfamiliar with formal verification.
4.2. Results
To understand if participants knew the icon’s (see Figure 2) meaning, we asked them to explain it. Of the 10 participants, 5 mentioned formal verification, and the other five mentioned concepts related to the security of the Password Manager (e.g., “the icon means that the passwords were safe” (P1)).
Some users correctly identified the formally verified features.
Another aspect we explore is whether participants understand which features are formally verified. Specifically, we found that: (a) 60% identified the generator and the storage as formally verified; (b) 30% stated that the whole Password Manager was formally verified; when asked why one user stated “I saw the icon in several places” (P2); (c) and, one user (10%) could not explain, stating they did not know.
Users may be more willing to use a formally verified Password Manager.
When comparing users’ responses regarding trust in a formally verified Password Manager vs one that is not formally verified, users stated trusting more the formally verified Password Manager.
4.2.1. Limitations and Motivation for Second User Study.
A key limitation is the small sample size, which may not reflect a diverse population. Furthermore, our explanation of formal verification may have led participants to confuse security and verification. To avoid that issue in our main study, we completely refrain from explaining formal verification; instead, we communicate about the guarantees it provides. Another limitation is using a specific Password Manager, which might threaten the study’s generality. We argue that it did not influence the participants’ responses significantly as the interface does not appear remarkably different from other browser extension Password Managers (e.g., LastPass (lastpass, ) or 1Password (1password, )). However, we still decided not to use any specific Password Manager interface in our main study to facilitate generalization. After gathering the main insights from this explorative study, we designed and deployed a large-scale quantitative study informed by this formative study.
5. Main Study
To address all the research questions mentioned in Section 1 we designed and conducted a 200-participant online survey to understand users’ perceptions of formal verification.
Table 1. Percentage of participants that agreed or strongly agreed that the factor would impact their willingness to use a PM.
Results Factors
69.0% being inexpensive
73.5% having support materials (e.g., tutorials)
76.5% being certified by Password Manager Security Group
78.0% being free
81.5% being made by a trustworthy and familiar company
86.5% being mathematically correct
87.0% being easy to use for first-time/beginner users
5.1. Structure
First Section.
One of our goals was to understand if formal verification in a Password Manager affects users’ willingness to use it. To understand this, we first had to make sure that users knew what a Password Manager was, as previous studies have shown that some users do not use Password Managers because they do not know that these tools exist (pearman2019people, ; stobert2014password, ; alkaldi2016people, ). To explain what a Password Manager is and to situate the participants in our survey, we first present them with a vignette scenario where they read a news story about what a Password Manager is and the role of the primary password.
After viewing this scenario, we show participants a survey question addressing RQ1. We ask “Imagine that you were thinking about using a Password Manager. What would make you more willing to use the Password Manager?” We then present participants with different factors to which they respond using a Likert (lazar2017research, ) 5-point agreement scale. Among these factors, we include formal verification. The reason we added other factors in addition to formal verification was to: (1) Prevent participants from realizing that the focus of the study was formal verification. We did this to prevent common biases where the users are inclined to agree with the researcher (merrett2006reflections, ); (2) Understand if users’ preference for formally verified Password Managers differed significantly from other factors. These factors were, for example, that the Password Manager provides support materials such as tutorials and is free (see Table 1). One of the factors used was a certification by the “Password Manager Security Group”. This factor did not represent an actual entity. It was added to understand if participants were inclined to agree that any factor was impactful without understanding or because it sounded good.
Finally, we omit the term ”formal verification” by using analogies when possible. The specific analogy we chose is related to something most participants had to deal with in their education — mathematics. Instead of saying that the Password Manager is formally verified, we state that it is ”mathematically correct, that is, its features are as trustworthy as a mathematical proof.” We do this to follow best practices (redmiles2017summary, ) and exclude jargon from our survey as most users are unfamiliar with formal verification. Additionally, media outlets have previously used mathematical analogies to explain formal verification to broader audiences, such as the BBC (rubens_2015, ) and Quanta Magazine (hartnett_2020, ).
Second Section.
After asking participants if they found that formal verification (among other factors) would impact their willingness to use a Password Manager, we asked them to elaborate. We presented participants with their answers and asked, respectively, ”In the previous question, you stated you felt more/less willing to use a Password Manager that is mathematically correct (…). Please state your reasons.”. We hoped to understand why participants valued (or not) formal verification in a Password Manager and what they associated it with. We also hoped to gauge if there was any misconception about what formal verification was for participants.
Third Section.
To answer RQ2 and RQ3, we gather common Password Manager features (e.g., Password Generator and Clipboard clearing). For each of these, we present scenarios representing the impact formal verification can have on each feature. For example, for the Password Generator, the policy compliance scenario is: ”Imagine that you are creating a new account on a website (e.g., Twitter, Facebook). To increase security, you ask the Password Manager to generate a password with seven characters and at least two numbers. However, the password generated does not include any numbers.”. After each scenario, and using a 5-point Likert (lazar2017research, ) scale, we ask users if that scenario would make them stop using a Password Manager. A feature may have multiple scenarios if formal verification can impact it differently. An example is, again, the password generator, where formal verification may help with policy compliance but also with the guarantee that the generator is truly unpredictable (i.e., all passwords have the same probability). The scenarios were developed iteratively and followed guidelines for vignette scenarios design (barter1999use, ). In total, we designed seven unique scenarios.
5.2. Recruitment
We recruited participants through Prolific. [^5] Prolific is a crowd-sourcing platform that enables large-scale user studies by connecting research and users https://prolific.com. We recruited 200 participants who were paid 2.65 GBP per submission (with one submission per user and an average time between 8 and 9 minutes). We deployed the with LimeSurvey. [^6] LimeSurvey is an open-source online survey tool https://www.limesurvey.org/.
Participant demographics
About half of our sample identified as male and half as female. Our sample skewed younger than the general population as most participants were less than 34 years (87%), and the average age was 26. Most participants knew what a Password Manager was before the study (77.5%), but a slight majority of them stated that they did not use one currently (52%). The rest of the users were divided between those who said they were currently using a Password Manager (40%), those who had used one in the past but not anymore (3.5%), and those who did not know (3.5%). The three Password Managers that participants mentioned the most were Google Chrome Password Manager (41%), Bitwarden (15%), and Apple Keychain (14%).
Nonetheless, due to a lack of understanding about what a Password Manager is some participants could answer that they had never used a Password Manager when they used a built-in Password Manager (e.g., when they received prompts to save a password in a browser). So, to understand the actual number of participants that use a Password Manager, we also asked participants if they saved their passwords in built-in Password Managers. As stated before, 52% of participants claimed they had never used a Password Manager before. However, of that 52%, 67% admitted to saving passwords when prompted (i.e., using built-in Password Managers). These results seem to imply that users use Password Managers without realizing they are using them. Reasons for these results could be a lack of technical knowledge, low computer literacy, and a lack of understanding about a Password Manager. We do not know if using a Password Manager without knowing impacts users’ perception of these tools. In total, 77% of participants used Password Managers to store and manage passwords.
5.3. Analysis
To understand if there was a significant difference between different survey questions, we used non-parametrical statistical tests as recommended by Lazar et al. (lazar2017research, ). We used Friedman’s ANOVA and Wilcoxon Signed-Rank Test with Bonferroni continuity correction (for repeated measures (lazar2017research, )). Non-nonparametric tests were used in previous studies on user perceptions of security issues (previous-usenix-nonparametrical, ).
We coded the open answers from the survey using an inductive coding strategy. Two coders created the codebook iteratively using inductive and hierarchical coding (lazar2017research, ). In total, two coders coded 200 answers. We used Cohen’s Kappa to ensure inter-rater reliability and finished the coding process with a Cohen’s Kappa of 0.81, which denotes a very high level of agreement. An interpretation of Cohen’s Kappa is that a value above 0.60 indicates “satisfactory reliability” and one above 0.80 ”near-perfect agreement”  (lazar2017research, ). All answers were coded individually. However, regular meetings happened to discuss ambiguities and possible changes to the codebook. The final codebook consisted of 13 codes (see Appendix H), and each participant’s answer could have multiple codes. We used a hierarchical coding scheme for some of the codes, for example, the code “Extra Security” had as sub-codes: “Extra security: password generation”, for when participants mentioned that the Password Manager could create strong passwords; “Extra security: secure storage”, for when the participant stated that their passwords were safer inside the Password Managers’ storage; and ”Extra security: Protection from unwanted access”, for when they mentioned that the Password Manager prevents third-parties from accessing their passwords. We also included more general codes for when participants were not specific in their answers or showed a lack of understanding about what was being asked (”Answer general to all Password Managers and not specific to formally verified Password Managers” and ”Just seems better”) and one code for answers that were unusable to us (”Answer not relevant”). Two answers were coded as unusable, leaving this study with 198 usable participant answers for this question.
Figure 3. Frequency of codes representing the reasons participants gave for ”agreeing” or ”disagreeing” that they felt willing to use a mathematically correct Password Manager.
5.4. Results
In this section, we address each RQ and its respective study insights.
5.4.1. Formal Verification and Willingness to use a Password Manager (RQ1)
In the survey’s first question, we asked participants whether formal verification impacted their willingness to use a Password Manager. To prevent biases, we compared seven positive factors (see Table 1). The factors are all things that could be perceived as desirable to participants, so it is no surprise that most participants ”agreed” or ”strongly agreed” would be more likely to use a Password Manager with any of those factors. What we want to understand is if their opinion is the same across all factors or if there are some factors that users value more. A Friedman test showed that there was a statistically significant difference in participants’ answers between the different Password Manager’s features with a $\chi^{2}(2)=51.42$ and a $p<0.05$. However, the following pairwise comparison of users’ answers using the Wilcoxon rank sum testdid not find a significant difference between most factors ($p>0.05$).
Comparing all the factors, we found that more than 86% of participants considered the use of formal verification (as described in Section 5.1) and being easy to use as important (they ”agreed” or ”strongly agreed” that these factors would make them more willing to use a Password Manager). The factors that users considered less critical (while still being essential for most) were being inexpensive (i.e., it costs money even if it is not expensive) and having support features (e.g., tutorials and help pages) with 31% and 27%, respectively, not thinking they would be more willing to use a Password Manager with them. This information can be found in Table 1. Our results suggest that users are more willing to use a formally verified Password Manager than a non-verified one. With this information in mind, we now aim to understand why participants answered our survey the way they did.
Understanding User’s Reasons.
This section analyzes the second survey section and aims to understand users’ reasons for varying formal verification. We separated the coding into the three questions that could be asked to participants — participants that valued formal verification, had no opinion, and found it undesirable (green, grey, and red in Figure 3). Additionally, depending on the size of the participant’s answer, one or more codes could be applied.
Most participants who valued formal verification in their Password Managers mentioned the use of mathematics as a reason why (67 participants (35%)). Participants seem to associate mathematics with a “logical” and certain behavior. One even stated that they preferred to use a formally verified Password Manager “because mathematics doesn’t lie” (P89). Another stated, “I like to use such products or services that have scientific proof of their effectiveness” (P147).
“Security” codes were also frequent (see Section 5.3). Some participants mentioned they believe formal verification added security (e.g., ”I think it would be more secure.” (P91)), and others gave more detailed explanations (e.g., “The main reason to start using a password manager is (…) the passwords generated are the strongest (…)” (P5)). Over 50% of the participants mentioned security as a reason why they valued formally verified Password Managers, with some mentioning more than one specific subcode of security in their answers (e.g., one participant stated that they preferred a formally verified Password Manager because ”…it will be more difficult to get hacked since it will create more secure passwords” (P162)).
Overall, participants mostly agreed that formal verification was something to be desired in a Password Manager. They provided several reasons, but the more frequent ones were related to security and the use of mathematics. On the other hand, participants who did not find formal verification relevant had a general distrust of Password Managers, found that formal verification was not essential for them, or simply did not understand the concept. For example, one of the participants directly mentioned, ”I just don’t believe in that kind of app. ” (P22). Lack of understanding was also a frequent code present in the answers of participants who had a neutral response to this survey question (see the bars in yellow in Figure 3). Lack of understanding of technical concepts has been the cause of users’ trust issues with other security software in the past (e.g., in the domain of cryptocurrencies, lack of knowledge on the subject may cause users to question the security of these systems (presthus2017motivations, )). Nonetheless, most participants who agreed that formal verification was important to them stated that security had something to do with it.
RQ1. How does formal verification impact users’ willingness to use Password Managers? Users seem to be more willing to use a formally verified Password Manager than a non-formally verified Password Manager; Our results also suggest that formal verification has positively affected some users’ trust; The main reasons users value formal verification are the use of mathematics and the extra security formal methods provide.
Table 2. P-values of pairwise comparisons of users’ answers in the scenarios question using the Wilcoxon rank sum test and Bonferroni continuity correction. See Table 4 for the scenario descriptions.
Scenario 1 Scenario 2 Scenario 3 Scenario 4 Scenario 5 Scenario 6 Scenario 7
Scenario 2 2e-16 - - - - - -
Scenario 3 0.00013 2e-16 - - - - -
Scenario 4 0.47137 2e-16 1.2e-05 - - - -
Scenario 5 0.00011 3.3e-10 3.2e-13 0.00220 - - -
Scenario 6 3.2e-05 1.3e-10 1.6e-14 0.00097 0.89316 - -
Scenario 7 6.4e-10 2.8e-06 2e-16 2.2e-07 0.03915 0.03915 -
Scenario 8 1.1e-14 0.00138 2e-16 1.8e-11 0.00024 0.00019 0.08562
5.4.2. Formally Verified Features (RQ2)
After learning that formal verification seems to impact users’ interest in using Password Managers, we are interested in learning their priorities, thus addressing RQ2. We realize that users may not have the expertise to judge what is essential and what should be verified in a Password Manager. However, our goal is to understand what features, if verified, would be more impactful for users. We want to understand how to meet users’ wants and maximize the impact of formal verification. A Friedman’s test showed that there was a statistically significant difference in participants’ answers between the different scenarios with a $\chi^{2}(2)=331.08$ and a $p<0.05$. The results of a pairwise comparison of users’ answers using the Wilcoxon rank sum test and Bonferroni continuity correction can be seen in Table 2. Our results suggest that most of the differences between scenarios are statistically relevant ($p<0.05$).
Table 3. Percentage of participants that agreed or strongly agreed the scenario would make them stop using a PM. S# corresponds to Scenario # (see Table 4 in the Appendix for scenario descriptions).
Results Scenarios
54.50% S2 Policy Compliance
70.50% S8 Ransomware/ Deleting your vault
76.00% S7 Synchronization
79.00% S5 Autofill
83.00% S6 Clipboard Clearing
86.50% S4 Primary Password Exposure
92.50% S1 Unpredictability
96.00% S3 Vault Exposure
Overall, participants seem to consider some features more critical than others. Our results suggest that the secure vault and its encryption are the most important – 96% of participants stated they agreed or strongly agreed that they would stop using a Password Manager if the respective scenario (S3) occurred. In order the most important features to verify seem to be the secure vault (S3), password generator (S1 and S2), login in the Password Manager (i.e., primary password security in S4), clipboard (S6), synchronization across devices (S7 and S8) – see Table 3.
RQ2. What features would users like to see formally verified in a Password Manager? The password vault’s security seems to be the feature that should be prioritized in a future implementation of a formally verified Password Manager; Surprisingly, verifying that the Password Manager’s passwords will not be lost does not seem to be important for users, as scenarios, where there was no third-party access to the password, were not as impactful; Ideally, all the features mentioned in this section should be formally verified, as over 50% of participants found their scenarios impactful.
5.4.3. Formal Verification Guarantees (RQ3)
The scenarios’ question also provides insights on whether participants value the guarantees of formal verification in Password Managers, thus answering RQ3.
As stated before, most scenarios were impactful for users, so in general, our results suggest that participants value the guarantees that formal verification provides. Additionally, we can divide our scenarios into the following categories:
  • Scenarios where a third-party learns all the passwords: scenarios ”Vault Exposure” (S3) and ”Primary Password Exposure” (S4);
  • Scenarios where a third party, over time, can learn a large number of passwords: scenarios ”Unpredictability” (S1), ”Autofill” (S5), and ”Clipboard Clearing” (S6);
  • Scenarios where the user loses access to all the passwords stored in the Password Manager’s vault: scenarios ”Synchronization” (S7) and ”Deleting your vault” (S8);
  • A scenario where the generator does not behave as intended and the passwords generated are not compliant with the users’ inputs: scenario ”Policy Compliance” (S2).
Two of the three most impactful scenarios described a situation where a third party could learn all the users’ passwords (S3 and S4). In these scenarios, 86.5% and 92.5% of participants stated they ”agreed” or ”strongly agreed” that they would stop using a Password Manager if it happened. These scenarios were related to the absence of formal verification where its guarantees would prevent the exposure of passwords to a third party. However, the second, fourth, and fifth most impactful scenarios (S1, S5, and S6) described a situation where a third party, over time, could learn several users’ passwords. In these, over 79% of participants ”agreed” or ”strongly agreed” that they would stop using a Password Manager if they happened. The other type of scenario, where the user loses access to all the passwords they have stored in the Password Manager’s vault (S7 and S8), was not as impactful for participants as the ones mentioned before. Finally, the last type of scenario described a situation where the Password Manager did not work as intended (S2), but its malfunction did not lead to a loss of passwords. It just inconvenienced users. This was the scenario that fewer participants found impactful. Interestingly, the least important feature in our hierarchy—the unpredictability of the password generator (S1)—is a feature that previous work has made efforts to formally verify (grilo22pwdgen, ).
Our results thus suggest that users value the guarantees that formal verification provides. Moreover, users value the guarantees more depending on the impact that said guarantees have. Guaranteeing that passwords are not leaked to third parties (e.g., by ensuring the cryptographic algorithms are well implemented) seems to be very important for participants.
RQ3: Do users value the guarantees that formal verification can provide in Password Managers? Our results suggest that users value the guarantees that formal verification can provide in Password Managers; Our data also seems to indicate that some guarantees are more important than others. For example, guaranteeing that the passwords are not leaked to third parties seems more relevant for participants than guaranteeing that they are synced in the cloud.
6. Discussion
In this section, we situate our results within the broader context of formal verification application. We discuss the main insights and suggest recommendations for future verification efforts and research. Our main insights include:
User Awareness and Perception.
An insight we got was that there is an overall low level of awareness among users about what formal verification is. Despite this, users react positively to formal verification in Password Managers. There appears to be a gap between the technical understanding of formal verification and the perceived value it adds to Password Managers. Some of our participants liked formal verification due to an inherent trust in the mathematical and technical rigor it implies, suggesting that even a superficial understanding or awareness of formal verification may influence user preferences.
User features prioritization.
A significant insight from this study is that users perceive some Password Manager features as more important. The security of the password vault and the reliability of password generation were highlighted as particularly critical features for users. This insight underscores the need for developers to prioritize formal verification efforts on features that most directly impact user trust and perceived security. Understanding user priorities can guide the allocation of development resources towards the aspects of Password Managers that users value more and enhance the effectiveness and attractiveness of formally verified Password Managers.
6.1. Recommendations
One of our goals was to identify how to increase the impact of formal verification in software. To achieve this, and based on the results of our user studies, we present several recommendations for developers of formally verified products.
1. Increase user familiarity with formal verification.
Our results suggest that users do not know what formal verification is. With this in mind, developers of formally verified software should try to convey to users the role of formal verification in their software. We suggest that these efforts should be focused on the specific consequences of formal verification in software instead of theoretical concepts. Avoiding technical language may facilitate the application of best practices such as using metaphors and avoiding jargon (redmiles2017summary, ).
Increasing this transparency may enable users to understand the advantages of formally verified software and, thus, increase its adoption. If formally verified software is valued by users, it may also increase its investment value for businesses (for example, for Password Managers such as Bitwarden (bitwarden, )).
2. Don’t overstate/mislead users when explaining formally verified software.
As mentioned before, some users associated formal verification with security (see Section 4.2). While this sometimes can be true, security is not necessarily related to formal verification. As such, we suggest that future efforts to communicate about formal verification should take care to prevent overstating the impact it has. Misleading users can be counterproductive and lead to distrust.
3. Make an effort to understand the users’ priorities in formally verified software.
Even if you do not intend to make formal verification a selling point of the product, we suggest that it is important to understand what are the users’ priorities. Understanding users’ wants can provide relevant insight into what can make a product more desirable. Moreover, since formally verifying software is time-consuming and expensive, understanding users’ priorities may help focus verification efforts on the features that are more impactful.
6.2. Limitations
User studies such as these may suffer from bias. Bias can arise from the questions or even the questionnaires. To mitigate our limitations, we performed cognitive interviews [^7] Cognitive interviews involve asking respondents to think aloud as they complete a survey and asking them questions about each survey item (redmiles2017summary, ). to validate the study until no more errors or typos were detected (we did two cognitive interviews for each study). We also followed best practices by offering “don’t know (DK)” or “prefer not to answer” responses (redmiles2017summary, ). Moreover, we asked users to explain their understanding of the topic to mitigate the Dunning-Kruger Effect (e.g., when they stated that they understand what formal verification is, we asked them to explain what it is) and included attention check questions. For example, we include one of these questions among the factors that may impact users’ willingness to use a Password Manager (see Section 5.1) and another among the scenarios in the third section of the survey (see Section 5.1).
We also removed as much jargon as possible. The jargon includes terms like “memory” and “encrypted” but also “formal verification” itself (redmiles2017summary, ) and tried to mitigate the Hawthorne effect [^8] Under the Hawthorne effect users are inclined to agree with researchers (merrett2006reflections, ). by hiding that the study is about formal verification. To prevent bias we randomized the order in which the factors that may impact users’ willingness to use a Password Manager are shown (see Section 5.1), and the order of the scenarios in the third survey section (see Section 5.1).
6.3. Future Work
Exploring formal verification’s impact on users’ willingness to use Password Managers opens up several avenues for further research.
Exploration in Other Domains.
Our study within the context of Password Managers suggests a positive user reception towards formal verification. Extending this research to other domains where formal verification is applied could offer an insightful research opportunity. For example, investigating user perceptions in the context of autonomous vehicles, medical devices, and blockchain technologies—all of which rely on the integrity and security assurances that formal verification provides—could reveal domain-specific user attitudes and expectations. These studies could help understand whether the positive inclination towards formal verification observed in Password Manager users is universal or if adjustments in communication and implementation strategies are needed based on the application domain.
Development of New Ways to Communicate about Formal Verification.
One significant finding from our study is the general lack of user awareness about formal verification. This gap presents a research opportunity to develop and test new communication methods about formal verification to the public. Inspired by initiatives like the IoT security labels(emami2020ask, ), a “Formal Verification Label” could be designed to provide at-a-glance information about the verification status of a product. Such a label could include simplified symbols or ratings that indicate the extent and areas of formal verification applied, making it easier for users to understand the claims of formally verified software products.
7. Conclusion
Formal software verification can be expensive and time-intensive, but our work suggests that it may positively impact users.This insight can be a powerful motivation for future formal verification efforts. In our study, we propose several directions for future research while providing concrete insights into which features should be prioritized by practitioners, such as enhancing the security of the password vault and ensuring the reliability of password generation. By focusing on these priorities, practitioners can allocate resources more efficiently, ensuring their efforts directly address the aspects of Password Managers that users value.Moreover, our work has shed light on a previously mostly unexplored area of research— combining formal verification with usable security. We argue that many paths for future work could be explored: virtually any domain where formal methods could be applied has to gain from studying user perception on the subject.
8. Acknowledgments
This work was funded by Fundação para a Ciência e a Tecnologia (FCT) under grant PRT/BD/153739/2021, and projects UIDB/50021/2020 (DOI: 10.54499/UIDB/50021/2020), LA/P/0063/2020 (DOI: 10.54499/LA
/P/0063/2020), and the PassCert project, a CMU Portugal Exploratory Project with reference CMU/TIC/0006/2019.
Appendix A Design of the Formal Verification Icon
Design of the Formal Verification Icon
The icon for formal verification should match the interface of Bitwarden, and because of that, we used the same font to design it (Font Awesome 4999Font Awesome is fully open source and GPL font friendly. https://fontawesome.com). Color is also very important: one of its roles is to contribute to the aesthetic value of the interface, but it can also be used to transmit information and influence the user emotionally (shneiderman2016designing, ). Although it is difficult to assume any universal interpretation of color, green is associated with safety in the Anglo-American cultural color convention (dix2004human, ). So, to differentiate from Bitwarden’s interface and draw attention to the formal verification logo, we used a new color, specifically green.
Bitwarden’s icons use two main colors, Bitwarden’s blue (#175DDC) and grey (#7C7C7C). Using a quadratic schema from Bitwarden’s blue and with the help of the Adobe Color [^10] A palette tool referenced by Tidwell et al. (tidwell2010designing, ). https://color.adobe.com/create, we derived the shade of green used in the icon (#0BDB0B).
It is also important to consider the needs of color-blind individuals. They may confuse some shades of orange or red with green or not see a red dot on a black background (shneiderman2016designing, ). Currently, in the Bitwarden interface, we do not have any colors that could be mistaken by color-blind users. Although we use green for the Bitwarden symbol, it remains legible (it is not a green text in a red background, for example).
The development of the icon went through three phases: (1) Together with a small group of colleagues developing research in security and formal methods, we began by brainstorming icon designs and ideas. After reaching 15 different variations, we had a group meeting to gather feedback. At this stage, we shortlisted 7 icons. (2) We then placed the 7 icon variations in their context (in the Password Managers interface), met again with the group, and asked for feedback. The team gave specific feedback about icon size, subjective aesthetic value, icon position in the tool, places where the icon should be, and lastly, about what icons represented more accurately formal verification. (3) To gather more unbiased feedback, we asked 20 users outside the team to give their opinions. Because the icon is intended to mean that a certain feature is formally verified, it is very important to the underlying message we want to transmit. As such, we performed user studies to determine the best variation of the icon. The participants were recruited through the authors’ personal network and did not receive compensation. The form where we asked for feedback about the icons was divided into three parts: (a) First, we asked users to choose the icons they liked more without knowing what they were meant to represent; (b) Secondly, we did a preference test, where we explained what the icon is trying to convey and asked users to again rate the icons according to the ones they prefer (harley_2016, ); (c) And lastly, a preference test, where we showed the icons in the context they would be in the final interface and asked users to rate them. This survey can be viewed in Appendix B, and the icon chosen can be seen in Figure 2. After an internal discussion about whether the icon could be confused with the HTTPS padlock, we argue that a padlock is not uniquely linked with the HTTPS security icon (von2022builds, ). Several icons were developed, and the final design was informed by several feedback meetings with the research team and a thorough user test, as described.
Appendix B Icon Survey
The survey used in the icons’ user study is divided in three main parts: Attractiveness Test (§B.2), Preference Test with no context (§B.3), and Preference Test with context (§B.4). The survey starts with the participants going through the informed consent form (§B.1).
B.1. Consent
The goal of this survey is to collect feedback on a set of icons that are developed to improve the user interface of password managers. The icons aim to convey that a certain component is formally verified (this is explained below).
This project is part of the PROJECT_NAME research project.
All the data collected is anonymous and will be used solely by the researchers of PROJECT_NAME. The data may be used to present insights at conferences, academic events, or similar events and for scientific publications.
Your participation is voluntary, and you may always quit at any time, without any kind of penalization. By selecting “Yes” below, you are consenting for your data to be processed, stored, and used as described above. You also are confirming that you have read this consent form.
The session is predicted to take about 4 minutes.
B.2. Part 1 – Attractiveness Test
Please order these icons from the one you find more aesthetically pleasing to the one you find least aesthetically pleasing (for example: E-G-F-H-I-K-J)
Figure 4. Icon options
B.3. Part 2 – Preference Test without context
Formal verification is a process in which developers prove (mathematically) that a certain part of a program behaves as intended. To say that a feature is formally verified means that it is guaranteed to behave a certain way.
A concrete example is guaranteeing that the length of generated passwords satisfies the minimum requirements.
Order these icons from the one you find that represents best the concept ”Formally Verified Feature” to the one that represents it worst (for example: E-G-F-H-I-K-J)
Figure 5. Icon options
B.4. Part 3 – Preference Test with context
Formal verification is a process in which developers prove (mathematically) that a certain part of a program behaves as intended. The icons bellow are meant to represent this concept.
1.
Considering the previous information, please order these images from the ones you like more and that represent the concept better to the worst (for example: E-G-F-H-I-K-J)
Figure 6. Icon options in the PM’s interface
2.
Considering the previous information, please order these images from the ones you like more and that represent the concept better to the worst (for example: E-G-F-H-I-K-J)
Figure 7. Icon options in the PM’s interface
3.
Considering the previous information, please order these images from the ones you like more and that represent the concept better to the worst (for example: E-G-F-H-I-K-J)
Figure 8. Icon options in the PM’s interface
4.
Thank you for filling this survey. Please use this space to give any feedback you consider relevant
Appendix C First User Study: Interview Protocol
This appendix shows the interview protocol followed for the first user study. Each of the nine steps includes the script followed.
1.
Introduce yourself. Give them the consent form to sign, on their worksheet. If it is not signed, do not proceed.
My name is NAME, and I will be giving you instructions on what to do and will answer your questions.
In this questionnaire the participant should read and accept the following consent form:
“The goal of this survey is to information about you, your experience with Password Managers and specifically with PROJECT_NAME’s Password Manager.
This project is part of the PROJECT_NAME research project, a project that is building an open-source, proof-of-concept password manager that through the use of formal verification, is guaranteed to satisfy properties on data storage and password generation.
All the data collected is anonymous and will be used solely by the researchers of PROJECT_NAME. The data may be used to present insights at conferences, academic events, or similar events and for scientific publications.
Your participation is voluntary, and you may always quit at any time, without any kind of penalty. By selecting ”Yes” below, you are consenting for your data to be processed, stored, and used as described above. You also are confirming that you have read this consent form.”
2.
Describe the purpose of the study.
The goal of this project is to evaluate the interface of PROJECT_NAME’s Password Manager. A password manager is a program used to store and manage passwords. This is a product in development and some formally verified features are not fully implemented. We’re looking for places where the product may be difficult to use. If you have trouble with some of the tasks, it’s the product’s fault, not yours. Don’t feel bad; that’s exactly what we’re looking for. If we can locate trouble spots, then we can go back and improve the product. Remember we’re testing the product, not you. If you have any feedback we ask you to be as honest as possible.
3.
Tell the participant that it’s OK to quit at any time.
Although I don’t know of any reason for this to happen, if you should become uncomfortable or find this test objectionable in any way, you are free to quit at any time.
4.
Explain how to think aloud
We have found that we get a great deal of information from these informal tests if we ask people to think aloud as they work through the exercises. It may be a bit awkward at first, but it’s really very easy once you get used to it. All you have to do is speak your thoughts as you work. If you forget to think aloud, I’ll remind you to keep talking. As you go through these tasks, we ask that you think aloud. You can say anything that comes to mind. This helps us better understand how users like yourself experience the product. If you forget to think aloud, I’ll remind you to keep talking.
5.
Explain that you will not provide help.
As you’re working through the tasks, I won’t be able to provide help or answer questions. This is because we want to create the most realistic situation possible. Even though I won’t be able to answer your questions, please ask them anyway. It’s very important that I capture all your questions and comments. When you’ve finished all the exercises, I’ll answer any questions you still have.
6.
Time and anonymity.
The results of this evaluation will be summarized and used in the context of the PROJECT_NAME research project. If the results of this experiment are published your identity will be anonymized.
The expected duration of this session is 50 minutes.
Do you have any questions at this point?
7.
Give the pre-study questionnaire.
Remember that a Password Manager is a software used to store and manage passwords. Before we begin, please fill out a pre-study questionnaire. This will provide us useful information about your demographics.
Before we start the tasks, I would like you to explore the product. Feel free to click any buttons and take your time.
8.
Go through the tasks in the interface. Ask participants to fill the Final Questionnaire.
Now I would like to ask you to fill the Final Questionnaire.
9.
Further feedback. Discuss any interesting behaviors you would like the participant to explain.
Do you have further feedback or suggestions regarding the product? Thank you for your participation. If you have any questions or wish to learn more about this research project do not hesitate to contact us.
Appendix D Pre Task Questionnaire
The pre-task questionnaire was used as the basis for a semi-structured interview where participants were encouraged to think-aloud about the questions. Most Likert Scale survey answers were accompanied by the participants reasons for their answers.
D.1. Part 1 – Password Managers
Q1.
Did you know what a password manager (PM) was before this study?
If you answered “yes” to the previous question please explain what a Password Manager is.
Q2.
Do you use a password manager?
Choose one of the following answers: Yes, I currently use a PM; Yes, I currently use a PM; No, I’ve never used one; I’m not sure.
Q3.
How do you manage/remember passwords?
D.2. Part 2 - Exclusively for participants that have never used a PM
Q1.
Before this study have you ever been interested in using a PM? Choose one of the following answers: Yes; No, and I don’t know what they are; No, but I know what Password Managers are.
Q2.
How much do you agree with the following statement:
I want to use a Password Manager in the future.
Choose one of the following answers: Strongly disagree, Disagree, Neither agree or disagree, Agree or Strongly Agree.
D.3. Part 3 - Exclusively for participants that are currently using a PM
Q1.
What Password Manager do you use?
Q2.
How often have you used your Password Manager (i.e. using the credentials in a Password Manager to log in websites)?
Q3.
In which context do you use a Password Manager (e.g. Work, Personal Life, etc)?
Q4.
What types of account do you save in the PM?
Choose all that apply: Social Media, Homebanking, Credit Cards, Online shopping, Mail, News/Entertainment, Other Options (open answer).
D.4. Part 4 - Exclusively for participants that have stopped using a PM
Q1.
What Password Manager did you use?
Q2.
Why did you stop using it?
Q3.
How much do you agree with the following statements:
I want to use a Password Manager in the future.
Choose one of the following answers: Strongly disagree, Disagree, Neither agree or disagree, Agree or Strongly Agree.
D.5. Part 5 – Background
Q1.
How much do you agree with the following statement:
  • I trust password managers (with my passwords).
  • I think Password Managers are, overall, difficult to use.
  • I feel my password are safer in a PM.
  • I feel Password Managers are, overall annoying to use.
  • I am willing to pay for software products.
  • I understand how Password Managers work.
Choose one of the following answers: Strongly disagree, Disagree, Neither agree or disagree, Agree or Strongly Agree.
Q2.
When using a browser do you allow the browser to remember your password?
Q3.
Do you know what formal verification is? Please explain if you do.
Q4.
How do you describe your gender identity?
Choose one of the following answers: Male; Female; Non-binary; Prefer not to say; Other.
Q5.
How old are you?
Choose one of the following answers: 18-24 years old; 25-34 years old; 35-44 years old; 45-54 years old; Above 55.
Q6.
What devices do you use daily?
Choose one, or more, of the following answers: Laptop; Desktop; Smartphone; Table; Other.
Q7.
What is the highest degree you have completed? (If currently enrolled, highest degree received.)
Choose one of the following answers: Less than a high school diploma High school degree or equivalent; Bachelor’s degree; Master’s degree; Doctorate degree.
Appendix E Final Questionnaire
The final questionnaire was used as the basis for a semi-structured interview where participants were encouraged to think-aloud about the questions. Most Likert Scale survey answers were accompanied by the participants’ reasons for their answers.
E.1. Part 1 – SUS
This section was composed of the standard SUS Questionnaire.
E.2. Part 2 – Knowledge about PMs
1.
What does this icon symbolize (in the context of PROJECT_NAME)?
See Figure 2.
2.
How much do you agree with the following statements:
  • I feel my password are less safe in a PM.
  • I feel I could trust in a Password Manager to save my passwords for me.
  • I know how a Password Manager works.
  • It is more convenient to use a Password Manager than to memorize passwords.
  • I think Password Managers are, overall, difficult to use.
  • If I were to use a Password Manager it would not be important for me to understand how the Password Manager works.
Choose one of the following answers: Strongly disagree, Disagree, Neither agree or disagree, Agree or Strongly Agree.
2.
When using a PM, how important is:
  • Ease of use.
  • Support material (tutorials, help pages).
  • Security and Privacy/Feeling that my data is secure.
  • Having formally verified features.
  • Price.
  • Understanding how the Password Manager works.
  • Being open source (i.e. the code is open for everyone to see).
  • The company that made and manages the PM.
Choose one of the following answers: Not at all important, Slightly Important, Important, Fairly Important or Very Important.
E.3. Part 2 – Impact of Formal Verification
1.
How much do you agree with the following statements:
  • I feel my password are less safe in a formally verified PM.
  • I feel I could trust in a formally verified Password Manager to save my passwords for me.
  • I know how a formally verified Password Manager works.
  • I prefer Password Managers that are not formally verified.
  • I would be willing to pay more for a software product if I knew it was formally verified.
  • I would not trust a product just because it was formally verified.
  • I do not value the use of Formal Verification in software products.
Choose one of the following answers: Strongly disagree, Disagree, Neither agree or disagree, Agree or Strongly Agree.
2.
Do you know what formal verification is? Please explain.
3.
In your understanding, what is formal verification?
4.
What specific features were formally verified in the Password Manager you used? What is the used of formal verification in the features you mentioned.
E.4. Part 3 – Perception of Safety and Final Questions
1.
After using PROJECT_NAME’s Password Manager how much do you agree with the following statements:
  • I felt safe using the PM.
  • I felt my password were secure in the PM.
  • I would be more willing to use a Password Manager in the future if I knew it was formally verified.
Choose one of the following answers: Strongly disagree, Disagree, Neither agree or disagree, Agree or Strongly Agree. Justify your answers.
2.
Will you use/continue to use a Password Manager in the future?
Appendix F Scenarios descriptions
Table 4. Scenarios descriptions.
PM Feature Name of Scenario Scenario
Password Generator Scenario 1 Unpredicta- bility Imagine that you use a Password Manager to generate several new random passwords for services that you use (e.g. Twitter, Facebook, internet banking). After generating a few passwords, you notice that all the generated passwords share a common pattern: they all start and end with the same set of letters. Because of this, the password was predicted by an attacker and now they have this password.
Scenario 2 Policy Compliance Imagine that you are creating a new account on a website (e.g. Twitter, Facebook). To increase security, you ask the Password Manager to generate a password with 7 characters and with at least 2 numbers. However, the password generated does not include any numbers.
Password Storage Scenario 3 Vault exposure Imagine that you are using a Password Manager to store your passwords and you discover that an attacker was able to access all your passwords saved in the Password Manager without knowing the primary password.
Primary Password Scenario 4 Primary Password Exposure Imagine that you log in to your Password Manager with your primary password, but it is exposed and, as a result, an attacker learns your primary password.
Autofill Scenario 5 Autofill The autofill feature of a Password Manager automatically fills your username and password in a website’s login page to make logging in easier. Imagine that you want to login in a website with a password that is saved in the Password Manager. The autofill feature fills in the password field and you log in. After this, you discover that an attacker was able to discover your password to that website because the autofill feature left the password exposed.
Clipboard Scenario 6 Clipboard Clearing Imagine that you want to log in to a website. To log in you decide to copy the password from the Password Manager and paste it into the website to log in. When you copy the password it can be pasted anywhere. But, to prevent an attack, 30 seconds after you copy the password, the Password Manager should clear it so an attacker can’t paste it again and discover it. However, this doesn’t happen and, as a result, an attacker learns your password.
Synchroniza- tion across devices Scenario 7 Synchroni- zation The Password Manager that you use allows you to keep a backup of all your passwords. Imagine now that an attacker deletes all your passwords, and when you try to restore them with the backup, you find that not all passwords were backed up and, as such, there are some passwords that you lost.
Scenario 8 Ransomware/ Deleting your vault The Password Manager that you use allows you to keep a backup of all your passwords. Imagine now that you lose access to your computer and need to access your passwords in the Password Manager from another location. However, you discover that not all passwords were backed up and, as such, there are some passwords that you lost.
Appendix G Second Study Survey
The survey used in the second study is divided in four main parts: Password Managers (§G.2), Mathematically Correct Password Manager (§G.3), Scenarios (§G.4), and Background (§G.5). The survey starts with the participants going through the informed consent form (§G.1).
G.1. Consent
Thank you for your interest in participating in the survey ”Users’ willingness to use Password Managers”. Below, you find information about this research project, conditions for participation, and handling of the collected data. Please read everything carefully. If you agree and wish to participate in this study, please confirm your consent below.
General information about the research project
This study is conducted in the context of the PROJECT_NAME. The goal of this survey is to investigate password managers and factors that impact their use.
In this study, you will be asked to complete a questionnaire about Password Managers. To this end, we will ask you to indicate how strongly you agree or disagree with statements and scenarios. We will ask you some demographic data. The survey will take approximately 10 minutes to complete. There are no specific prerequisites for participating in this study. Please note that you must be at least 18 years old to participate in this study. You can earn £1.65 for your participation. No particular burdens or damages from participating in this research project are to be expected.
Voluntariness
Your participation in this research project is voluntary. You can revoke your consent to participate at any time, without providing reasons, and without any disadvantages.
Responsible management of the research project
If you have any questions regarding the research project or if you wish to make use of your right of revocation, please contact:
NAME and EMAIL ADDRESS
There are 25 questions in this survey.
G.2. Part 1 – Password Managers
Please read the following scenario carefully.
Imagine that you saw a news story about the use of Password Managers. In the news, they stated that Password Managers are programs that store and generate passwords. They stated that Password Managers can help increase your password security online. With a Password Manager, you save all your passwords in a secure vault and only have to remember one password — the primary password. Imagine now that you were thinking about using a Password Manager.
How much do you agree with the following options:
I would be more willing to use a Password Manager that…
  • has support materials like tutorials or help pages
  • is mathematically correct, that is, its features are as trustworthy as a mathematical proof
  • is inexpensive
  • is free
  • is certified by PMSG (Password Manager Security Group)
  • is made by a trustworthy company that is familiar to you
  • please select the option “Strongly disagree”
  • is easy to use for first-time and beginner users
Please choose the appropriate response for each item: Strongly disagree, Disagree, Neither agree or disagree, Agree or Strongly agree.
G.3. Part 2 - Mathematically Correct Password Manager
Depending of the answer to the previous question there are three possible questions that could be asked in this part of the survey.
G.3.1. Option A
Only answer this question if answer was ‘Agree’ or ‘Strongly agree’ at question How much do you agree with the following options: I would be more willing to use a Password Manager that… (is mathematically correct, that is, its features are as trustworthy as a mathematical proof).
In the previous question you stated you felt more willing to use a Password Manager that is mathematically correct, that is, its features are as trustworthy as a mathematical proof. Please state your reasons.
G.3.2. Option B
Only answer this question if the following conditions are met: Answer was ‘Strongly disagree’ or ‘Disagree’ at question How much do you agree with the following options: I would be more willing to use a Password Manager that… (is mathematically correct, that is, its features are as trustworthy as a mathematical proof).
In the previous question, you stated you felt less willing to use a Password Manager that is mathematically correct, that is, its features are as trustworthy as a mathematical proof. Please state your reasons.
G.3.3. Option C
Only answer this question if the following conditions are met: Answer was ‘Neither agree or disagree’ at question How much do you agree with the following options: I would be more willing to use a Password Manager that… (is mathematically correct, that is, its features are as trustworthy as a mathematical proof).
In the previous question, we asked you if you felt more willing to use a Password Manager that is mathematically correct, that is, its features are as trustworthy as a mathematical proof. You stated that you neither agree nor disagree with this statement. Please state your reasons.
G.4. Part 3 – Scenarios
Remember that Password Managers are programs that store and generate passwords. With a Password Manager, you save all your passwords in a secure vault and only have to remember one password — the primary password.
In this section we’ll present you with scenarios and for each we ask how much do you agree with the following:
I would stop using a Password Manager if this scenario happened.
Please read each scenario carefully.
G.4.1. Scenario A
I would stop using a Password Manager if this scenario happened:
Imagine that you use a Password Manager to generate several new random passwords for services that you use (e.g. Twitter, Facebook, internet banking). After generating a few passwords, you notice that all the generated passwords share a common pattern: they all start and end with the same set of letters. Because of this, the password was predicted by an attacker and now they have this password.
Choose one of the following answers: Strongly disagree, Disagree, Neither agree or disagree, Agree or Strongly Agree.
G.4.2. Scenario B
I would stop using a Password Manager if this scenario happened:
Imagine that you are creating a new account on a website (e.g. Twitter, Facebook). To increase security, you ask the Password Manager to generate a password with 7 characters and with at least 2 numbers. However, the password generated does not include any numbers.
Choose one of the following answers: Strongly disagree, Disagree, Neither agree or disagree, Agree or Strongly Agree.
G.4.3. Scenario C
I would stop using a Password Manager if this scenario happened:
Imagine that you are using a Password Manager to store your passwords and you discover that an attacker was able to access all your passwords saved in the Password Manager without knowing the primary password.
Choose one of the following answers: Strongly disagree, Disagree, Neither agree or disagree, Agree or Strongly Agree.
G.4.4. Scenario D
I would stop using a Password Manager if this scenario happened:
Imagine that you log in in your Password Manager with your primary password, but it is exposed and, as a result, an attacker learns your primary password.
Choose one of the following answers: Strongly disagree, Disagree, Neither agree or disagree, Agree or Strongly Agree.
G.4.5. Scenario E
I would stop using a Password Manager if this scenario happened:
The autofill feature of a Password Manager automatically fills your username and password in a website’s login page to make logging in easier. Imagine that you want to log in in a website with a password that is saved in the Password Manager. The autofill feature fills in the password field and you log in. After this you discover that an attacker was able to discover your password to that website because the autofill feature left the password exposed.
Choose one of the following answers: Strongly disagree, Disagree, Neither agree or disagree, Agree or Strongly Agree.
G.4.6. Scenario F
I would stop using a Password Manager if this scenario happened:
Imagine that you want to log in to a website. To log in you decide to copy the password from the Password Manager and paste it into the website to log in. When you copy the password it can be pasted anywhere. But, to prevent an attack, 30 seconds after you copy the password, the Password Manager should clear it so an attacker can’t paste it again and discover it. However, this doesn’t happen and, as a result, an attacker learns your password.
Choose one of the following answers: Strongly disagree, Disagree, Neither agree or disagree, Agree or Strongly Agree.
G.4.7. Scenario G
I would stop using a Password Manager if this scenario happened:
The Password Manager that you use allows you to keep a backup of all your passwords. Imagine now that an attacker deletes all your passwords, and when you try to restore them with the backup, you find that not all passwords were backed up and, as such, there are some passwords that you lost.
Choose one of the following answers: Strongly disagree, Disagree, Neither agree or disagree, Agree or Strongly Agree.
G.4.8. Scenario H
I would stop using a Password Manager if this scenario happened:
The Password Manager that you use allows you to keep a backup of all your passwords. Imagine now that you lose access to your computer and need to access your passwords in the Password Manager from another location. However, you discover that not all passwords were backed up and, as such, there are some passwords that you lost.
Choose one of the following answers: Strongly disagree, Disagree, Neither agree or disagree, Agree or Strongly Agree.
G.4.9. Scenario I
I would stop using a Password Manager if this scenario happened:
To get results that are relevant, all participants need to pay attention to the scenarios in the survey. This question is to make sure you are paying attention to the study, so please choose the answer ”strongly agree“.
Choose one of the following answers: Strongly disagree, Disagree, Neither agree or disagree, Agree or Strongly Agree.
G.5. Part 4 – Background
1.
Are you familiar with the concept of formal verification?
Choose one of the following answers: Yes, No or I’m not sure
2.
If you are familiar with the concept of formal verification, in your understanding, what is formal verification?
3.
Did you know what a password manager was before this study?
Choose one of the following answers: Yes, No or I’m not sure
4.
Have you used a password manager in the past? Choose one of the following answers:
  • Yes, and I currently use one;
  • Yes, but not anymore;
  • No;
  • I don’t know
5.
Which password manager did you use?
Only answer this question if answer was ’Yes, and I currently use one’ or ’Yes, but not anymore’ at question Have you used a password manager in the past?.
6.
How do you remember your passwords? Select all that apply.
Check all that apply:
  • I save my passwords in a password manager
  • I write my passwords down on paper
  • I write my passwords in documents (e.g. text files or excel sheets)
  • I remember my passwords from memory
  • Other:
7.
When you use a login in a website sometimes a pop-up appears asking you to save the password for that website (this pop-up can appear, for example, in Chrome, Firefox, Safari, Apple Keychain). Do you save your passwords when prompted?
Choose one of the following answers:
  • Yes, I always save my passwords when asked;
  • Yes, I sometimes save my passwords when asked;
  • No, but I’ve saved passwords when asked in the past;
  • No, I never save my passwords when asked;
  • I don’t know
G.6. Part 5 - Demographics
1.
How do you describe your gender identity?
Choose one of the following answers: Male; Female; Non-binary; Prefer not to say; Other.
2.
How do you describe your race or ethnic identity?
Check all that apply: Black; White; Hispanic; Asian; Native American; Prefer not to say; Other.
Your survey responses have been recorded.
Thank you for completing this survey.
Appendix H Codebook used in the second study
Table 5. Codebook used in second user study
Code Description of code
Frequency
of Code
Extra security
The participants perceive the Password Manager more positively because it is more secure or with more safety. But they
do not specify in what way it is more secure.
47
Extra security: password
generation
The participants perceive the Password Manager more positively because it is more secure. They specify that the password
generation is more secure.
26
Extra security: protection
from unwanted access
The participants perceive the Password Manager more positively because it is more secure. They specify that it can prevent
attacks from bad third parties OR specify that it is harder to hack the password.
29
Extra security: secure
storage
The participants perceive the Password Manager more positively because its more secure. They specify that the storage is
more secure. For example there is low chance of password leaks
18
Use of mathematics The participant mentions that the Password Manager employ scientific or mathematically principles. 67
Just seems better The participant believes that a Password Manager with formal verification ”just seems better” with no specific reason. 8
Trustworthy
The participant thinks they would trust this Password Manager more and that is why they would like to use a mathematical
correct PM.
39
More assurances The behavior of the Password Manager is more certain. This may not have anything to do with security (e.g. less bugs). 14
Distrust of Password Managers The participant does not trust in Password Managers in general. 5
Not relevant The participant finds that even if a Password Manager is mathematically correct this is not relevant for them. 12
Lack of understanding
Lack of understanding about what mathematically correct means makes users not sure if it’s good OR about
what is being proposed in the question and survey.
30
Not related to verification
or mathematics.
The participant response does not seem related to verification or mathematics. 5
Answer not relevant The answer is not relevant to this study. 2

### v2

@@
 Are Users More Willing to Use Formally Verified Password Managers?
-Abstract.
+Abstract
 Formal verification has recently been increasingly used to prove the correctness and security of many applications.
@@
 However, while most security experts recognize the value of formal verification, the views of non-technical users on this topic are unknown.
-To address this issue, we designed and implemented two experiments to understand how formal verification impacts users.
+We designed and implemented two experiments to address this issue to understand how formal verification impacts users.
 Our approach started with a formative study involving 15 participants, followed by the main quantitative study with 200 individuals.
-We focus on the application domain of password managers since it has been documented that the lack of trust in password managers might lead to lower adoption.
-Moreover, recent efforts have focused on formally verifying (parts of) password managers.
-We conclude that formal verification is seen as desirable by users and identify three actional recommendations to improve formal verification communication efforts.
-Password Manager, Usable Security, Formal Methods, Formal Verification, User Study, HCI
-journal: JACMjournalvolume: 37journalnumber: 4article: 111publicationmonth: 8ccs: Security and privacy Usability in security and privacyccs: Security and privacy Formal methods and theory of security
-1. Introduction
+We focus on the application domain of Password Managers (PMs) since it has been documented that the lack of trust in PMs might lead to lower adoption.
+Moreover, recent efforts have focused on formally verifying (parts of) PMs.
+We conclude that formal verification is seen as desirable by users and identify three actionable recommendations to improve formal verification communication efforts.
+Keywords:
+Password Managers, Usable Security, Formal Methods, Formal Verification, User Study, HCI, User Trust, Security Adoption, Verified Software, Software Assurance, Human Factors in Security, Cryptography Usability, Secure Software Adoption
+1 Introduction
 Formal verification has developed substantially in recent years and has been applied to many domains.
-Recent work in this area includes the formal verification of security properties in trusted execution environments (jangid2021towards, ), critical parts of the Linux kernel (nelson2020specification, ), the open-source TLS implementation used in numerous Amazon services (chudnov2018continuous, ), and even compilers (leroy2006formal, ; kumar2014cakeml, ).
+Recent work in this area includes the formal verification of security properties in trusted execution environments [27], critical parts of the Linux kernel [40], the open-source TLS implementation used in numerous Amazon services [10], and even compilers [34, 31].
 Formal verification works by rigorously proving the correctness of code against a formal specification and thus ensuring that software components behave as intended, even under adversarial conditions.
-An extensive survey conducted in 2020 (2020_expert_surveyFM, ) revealed that experts in formal methods widely believe these approaches offer enhanced code quality, strengthened cybersecurity, streamlined certification processes, and more manageable maintenance efforts.
-Although security experts generally recognize the value of formal verification, the views (or even awareness) of end-users are unknown (carreira2021exploring, ; carreira2022studying, ).
-Previous research has explored user expectations and perceptions of security in areas such as differential privacy (cummings2021need, ) and cryptocurrency systems (mai2020user, ).
-Similarly, prior work on formal methods has also successfully evidenced trends, limitations, and future paths for formal methods (2020_expert_surveyFM, ).
+An extensive survey conducted in 2020 [18] revealed that experts in formal methods widely believe these approaches offer enhanced code quality, strengthened cybersecurity, and more manageable maintenance efforts.
+Although security experts generally recognize the value of formal verification, the views (or even awareness) of end-users are unknown [8, 6].
+Previous research has explored user expectations and perceptions of security in areas such as differential privacy [11] and cryptocurrency systems [37].
+Similarly, prior work on formal methods has also successfully evidenced trends, limitations, and future paths for formal methods [18].
 However, to the best of our knowledge, the impact of formal verification on end-user willingness to use formally verified products is still unexplored.
 Understanding users’ views on formal verification can help foster the adoption of formally verified code (with all the advantages verified code can provide) and aid developers in communicating with end-users and industry partners.
-Our work addresses this gap in the literature by studying formal verification’s impact on users of a concrete application domain: Password Managers.
-Password Managers are applications that help users generate and manage their passwords.
-Most Password Managers offer password security features like password generation, multi-factor authentication, and secure storage with a primary password.
-[^1] Some related work (pearman2019people, ; ray2020older, ; chiasson2006usability, ) uses “master password” instead of “primary password”.
-However, recent work uses “primary password” (e.g., Firefox’s Password Manager (mozilla, )).
-We chose this domain because text passwords are one of the most used security mechanisms, and Password Managers are an essential security tool to manage them (persistenceOfPasswords, ).
-However, previous work suggests that users have trust issues with Password Managers (ion2015no, ; ray2020older, ) and are reluctant to use them due to a lack of understanding of their security properties (pearman2019people, ).
-In parallel, recent work on formal verification has been applied to Password Managers (grilo2021towards, ; grilo22pwdgen, ).
+Our work addresses this gap in the literature by studying formal verification’s impact on users of a concrete application domain: Password Managers (PMs).
+PMs are applications that help users generate and manage their passwords.
+Most PMs offer password security features like password generation, multi-factor authentication, and secure storage with a primary password.
+We chose this domain because text passwords are one of the most used security mechanisms, and PMs are an essential security tool to manage them.
+However, previous work suggests that users have trust issues with PMs [26, 45] and are reluctant to use them due to a lack of understanding of their security properties [41].
+In parallel, recent work on formal verification has been applied to PMs [20].
 We designed and deployed two studies to understand users’ views on formal verification.
-The first is a formative interview study, asking users to use a Password Manager with verified components.
-The main study is a more extensive quantitative study that builds upon the formative and aims to explore some of the findings and themes identified in more detail.
+The first is a formative interview study.
+The main study is a more extensive quantitative study that builds upon the formative and aims to explore some of its findings in more detail.
 Contributions.
 Overall, our main contributions are:
-Addressing a critical gap in the literature: Our research provides an empirical evaluation of the influence of formal verification on end-users within a specific application domain —- Password Managers.
-This work makes a contribution by addressing a gap in the formal methods literature – the study of end-users of formal verification.
-Impact of formal verification on user adoption: Our findings suggest that formal verification has a positive effect on users’ willingness to adopt Password Managers.
-In both our formative user study and main study, participants showed an increased willingness to use Password Managers with formally verified components and identified formal verification as a desirable feature.
-Identification of features for formal verification: We systematically identify a hierarchy of Password Manager features that users consider most critical for formal verification.
+Addressing a critical gap in the literature: Our research provides an empirical evaluation of the influence of formal verification on end-users within a specific application domain — PMs.
+This work makes a contribution by addressing a gap in the formal methods literature — the study of end-users of formally verified products.
+Impact of formal verification on user adoption: Our findings suggest that formal verification has a positive effect on users’ willingness to adopt PMs.
+In both our formative user study and main study, participants showed an increased willingness to use PMs with formally verified components and identified formal verification as a desirable feature.
+Identification of features for formal verification: We systematically identify a hierarchy of PM features that users consider most critical for formal verification.
 We identify several critical features that are particularly relevant for participants.
@@
 We argue that practitioners can increase the appeal and effectiveness of verified security software by: i) addressing user-identified priorities, ii) enhancing transparency around the function and benefits of formal verification, and iii) engaging in broader educational initiatives to educate users about formal verification.
-2. Scope and Research Questions
-Our formative study motivates three questions.
-This study aims to test users’ perceptions of formal verification.
-To this end, we choose the domain of Password Managers and specifically address the following Research Questions:
-RQ1.:
-How does formal verification impact users’ willingness to use Password Managers?
-RQ2.:
-What features would users like to see formally verified in a Password Manager?
-RQ3.:
-Do users value the guarantees formal verification can provide in Password Managers?
+2 Scope and Research Questions
+Our study aims to test users’ perceptions of formal verification. To this end, we choose the domain of PMs and address the following Research Questions:
+RQ1.
+How does formal verification impact users’ willingness to use PMs?
+RQ2.
+What features would users like to see formally verified in a PM?
+RQ3.
+Do users value the guarantßees formal verification can provide in PMs?
 Summary of Methods.
-Figure 1. Overview of the Methodology
-Our overarching goal is to understand how formal verification in Password Managers influences the decisions and trust of non-expert users.
+Our overarching goal is to understand how formal verification in PMs influences the decisions and trust of non-expert users.
 The formative study, characterized by its qualitative and exploratory nature, focuses solely on RQ1.
-It aims to capture users’ initial reactions and understandings of formal verification in Password Managers through in-depth interviews and interactions with a Password Manager prototype.
-This prototype, an extended version of Bitwarden (bitwarden, ), incorporates a formal verification icon and explanations to educate users about formal verification’s role in enhancing Password Manager security.
-It gathers insights on general themes around Password Managers and formal verification, setting the groundwork for the subsequent larger-scale survey.
-The emphasis here is on identifying and exploring users’ perceptions, which directly contributes to addressing RQ1 by revealing how users view and understand formal verification in the context of Password Managers.
-The main study builds on the findings from the formative study, expanding the scope to address all research questions mentioned in the introduction.
-With 200 participants, this larger-scale survey aims to quantitatively assess the broader implications of formal verification on users’ willingness to use Password Managers.
-This study aims to validate the themes identified in the formative study at a larger scale and further explore how formal verification impacts users’ choices and trust in Password Managers.
-We use a combination of statistical analysis and scenario-based questions to understand better the role of formal verification across different aspects of Password Manager usage and answer our research questions.
-Crucially, we do not wish to ”explain” formal verification to our participants; we merely want to communicate its consequences and impact on the Password Manager.
-An overview of the methodology adopted can be seen in Figure 1.
+It aims to capture users’ initial reactions and understandings of formal verification in PMs through in-depth interviews and interactions with a PM prototype.
+This prototype, an extended version of Bitwarden [5] incorporates a formal verification icon and explanations to educate users about formal verification’s role in enhancing PM security.
+It gathers insights on general themes around PMs and formal verification, setting the groundwork for the subsequent larger-scale survey.
+The emphasis here is on identifying and exploring users’ perceptions, which directly contributes to addressing RQ1 by revealing how users view and understand formal verification in the context of PMs.
+The main study builds on the findings from the formative study, expanding the scope to address all RQs mentioned above.
+With 200 participants, this larger-scale survey aims to quantitatively assess the broader implications of formal verification on users’ willingness to use PMs.
+This study aims to validate the themes identified in the formative study at a larger scale and further explore how formal verification impacts users’ choices and trust in PMs.
+We use a combination of statistical analysis and scenario-based questions to better understand the role of formal verification across different aspects of PM usage and answer our RQs.
+Crucially, we do not wish to “explain” formal verification to our participants; we merely want to communicate its consequences and impact on the PM.
+An overview of the methodology adopted can be seen in Figure˜1.
 Ethical Considerations.
 In both studies, we did not collect any personal data.
-All participants were over 18 years old and were shown a consent form, the terms of which they had to accept before starting the survey.
+All participants were over 18 years old and were shown a consent form, which they had to accept before starting the survey.
 In the first author’s institution (the lead institution for this project), studies with these properties do not need to be submitted to the Ethics Committee (as confirmed by the Chair of the Ethics Committee).
@@
 All studies described in this paper followed these guidelines.
-Co-authors from other institutions did not have access to individual survey data and only worked with aggregate statistics, as reported in this paper.
-3. Background and Motivation
+Figure 1: Overview of the Methodology
+3 Background and Motivation
 Formal verification.
 In general, it is hard to build secure computer-based systems.
-Formal methods offer the promise of software that does not have exploitable bugs (fisher2017hacms, ) and has been used in a wide range of domains to prove the correctness and security of applications (jangid2021towards, ; nelson2020specification, ; grilo22pwdgen, ; chudnov2018continuous, ; ferreira2017certified, ; grilo2022verified, ).
+Formal methods offer the promise of software that does not have exploitable bugs and has been used in a wide range of domains to prove the correctness and security of applications [27, 40, 20, 10, 16].
 The idea of formal verification is mathematically modeling a system and then using formal methods to prove that the model satisfies specific properties or specifications.
 For example, Chudnov et al.
-(chudnov2018continuous, ) describe the development and operation of a continuously checked proof ensuring key properties of the TLS implementation used by many AWS services.
+[10] describe the development and operation of a continuously checked proof ensuring key properties of the TLS implementation used by many AWS services.
 Nelson et al.
-(nelson2020specification, ) describe their experience applying formal methods to a critical component in the Linux kernel, the just-in-time compilers.
+[40] describe their experience applying formal methods to a critical component in the Linux kernel, the just-in-time compilers.
 Their results show that building a verified component within a sizeable unverified system is possible with careful specification and proof strategy design.
@@
 Johnson et al.
-(johnson2020skeptic, ) introduce Skeptic, a toolchain that applies formal verification and power-law modeling to automatically select and justify password composition policies.
+[28] introduce Skeptic, a toolchain that applies formal verification and power-law modeling to automatically select and justify password composition policies.
 Formally verified compilers are also available.
-An example is the CompCert (leroy2009formal, ) compiler, which compiles code from a large subset of the C programming language to PowerPC assembly code and guarantees that safety properties proved on the source code hold for the executable compiled code as well (leroy2009formal, ).
-Password managers.
-Passwords have become increasingly complex as most websites adopt strict password security policies.Previous work suggests users struggle with generating unique passwords (inglesant2010true, ), and password reuse is a problem—Pearman et al.
-(pearman2017let, ) show that 40% of users reuse 81–90% of their passwords.
-The basic features that all Password Managers provide to users are: (a) A secure encrypted vault to save passwords; (b) A password generator to generate unpredictable passwords that comply with complex password security policies.
-Moreover, most modern Password Managers also provide other features, including: (c) “Autofill” to automatically fill passwords in a website’s login screen; (d) Cloud synchronization to keep passwords safe and synchronized across devices.
-Despite being recommended by experts  (enisapassword, ), Password Manager usage is not widespread.
-Users complain about not knowing that Password Managers exist and not understanding how they work (pearman2019people, ; ion2015no, ; ray2020older, ; chiasson2006usability, ).
-Users also do not trust Password Managers (pearman2019people, ).
-As an intermediary tool, Password Managers must be reliable, consistent, and predictable.
-Previous work (chiasson2006usability, ; pearman2019people, ) recommends education as a tool for enhancing adoption and trust in Password Managers.
+An example is the CompCert [35] compiler, which compiles code from a large subset of the C programming language to PowerPC assembly code and guarantees that safety properties proved on the source code hold for the executable compiled code as well [35].
+PMs.
+Passwords have become increasingly complex as most websites adopt strict password security policies.
+Previous work suggests users struggle with generating unique passwords [25], and password reuse is a problem — Pearman et al.
+[42] show that 40% of users reuse 81–90% of their passwords.
+The basic features that all PMs provide to users are: (a) A secure encrypted vault to save passwords; (b) A password generator to generate unpredictable passwords that comply with complex password security policies.
+Moreover, most modern PMs also provide other features, including: (c) “Autofill” to automatically fill passwords in a website’s login screen; (d) Cloud synchronization to keep passwords safe and synchronized across devices.
+Despite being recommended by experts  [14], PM usage is not widespread.
+Users complain about not knowing that PMs exist and not understanding how they work [41, 26, 45, 9].
+Users also do not trust PMs [41].
+As an intermediary tool, PMs must be reliable, consistent, and predictable.
+Previous work [9, 41] recommends education as a tool for enhancing adoption and trust in PMs.
 Usable security.
 An effective security mechanism is one that is used correctly.
-Usable security can be traced back to Saltzer and Schroeder’s 1974 paper (saltzer1975protection, ) that introduces the term “psychological acceptability” for access-control systems.
-In the last decade, there has been a lot of work on usable security across many domains (mai2020user, ; pearman2019people, ; ion2015no, ; carreira2021towards, ).
+Usable security can be traced back to Saltzer and Schroeder’s 1974 paper [48] that introduces the term “psychological acceptability” for access-control systems.
+In the last decade, there has been a lot of work on usable security across many domains [37, 41, 26, 7].
 Prior work explored various dimensions of security warnings, from SSL warnings in web browsers to end-to-end encryption guarantees.
 A notable contribution is the series of studies by Felt et al.
-(felt2015improving, ) examining the impact of SSL warnings on user behavior within Google Chrome.
+[15] examining the impact of SSL warnings on user behavior within Google Chrome.
 Their work focused on experimenting with SSL warnings and assessing user adherence and comprehension.
-Similarly, Akhawe and Felt’s (akhawe2013alice, ) study provided evidence on the effectiveness of browser security warnings, revealing insights into how users perceive and react to these alerts in real-world settings.
-Other relevant, usable security work includes previous efforts to communicate about security using labels similar to “nutrition labels (kelley2009nutrition, ).” In this realm, previous research designed privacy labels for several domains, such as app stores (li2022understanding, ) and IoT devices (emami2020ask, ).
-Related work has also evaluated the effectiveness of commonly deployed password strength meters and found significant inconsistencies between meter classifications and actual password resistance to cracking: many passwords labeled as “strong” were, in fact, easily guessed, and some “weak” passwords resisted cracking (pereira2020evaluating, ).
+Similarly, Akhawe and Felt’s [2] study provided evidence on the effectiveness of browser security warnings, revealing insights into how users perceive and react to these alerts in real-world settings.
+Other relevant, usable security work includes previous efforts to communicate about security using labels similar to “nutrition labels [30].” In this realm, previous research designed privacy labels for several domains, such as app stores [36] and IoT devices [13].
+Related work has also evaluated the effectiveness of commonly deployed password strength meters and found significant inconsistencies between meter classifications and actual password resistance to cracking: many passwords labeled as “strong” were, in fact, easily guessed, and some “weak” passwords resisted cracking [43].
 Limitations of Existing Approaches.
-A 2020 survey (2020_expert_surveyFM, ) of formal methods experts studied their views of formal methods.
-However, to the best of our knowledge, no previous work has focused on (non-expert) users of formally verified Password Managers, nor, more generally, on the impact that formal verification may have on end users.
-Previous work on formal verification of Password Managers is scarce, but efforts have been made to formally verify the random password generation algorithm of a Password Manager (grilo2021towards, ; grilo22pwdgen, ).
-However, these studies concentrated on the technical aspects of formal verification without examining its usability implications.
-4. Formative Study
+A 2020 survey [18] of formal methods experts studied their views of formal methods.
+However, to the best of our knowledge, no previous work has focused on (non-expert) users of formally verified PMs, nor, more generally, on the impact that formal verification may have on end users.
+Previous work on formal verification of PMs is scarce, but efforts have been made to formally verify the random password generation algorithm of a PM [20] or look at usability [6].
+However, these studies concentrated on the technical aspects of formal verification without examining its usability implications or had a smaller sample size.
+4 Formative Study
 Our goal with this formative study was to understand users’ overall views on formal verification.
-To achieve this, we implemented two significant extensions to a Password Manager: a formal verification icon and an explanation.
-4.1. Method
-We created a proof-of-concept prototype by extending an established Password Manager’s – Bitwarden’s (bitwarden, ) browser extension – interface and testing it for user acceptance.
+To achieve this, we implemented two significant extensions to a PM: a formal verification icon and an explanation.
+4.1 Method
+We created a proof-of-concept prototype by extending an established PM’s — Bitwarden’s browser extension — interface and testing it for user acceptance.
 Formal verification icon.
-Bitwarden uses icons to communicate with users.
-Similarly, we use an icon to communicate formal verification as icons are a good alternative to text to communicate with users (Wiedenbeck1999, ).
-The formal verification icon we chose is shown in Figure 2 and was placed on all features that could be formally verified: (a) Password vault, by the password field; (b) Primary password input box; (c) Password generator; (d) Clipboard, by its settings.
-We chose green due to its association with safety in Anglo-American contexts (dix2004human, ) and derived the green used in the UI from Bitwarden’s blue (#175DDC ) and grey (#7C7C7C ), ending up with the shade (#0BDB0B ).
-We designed the formal verification icon with Font Awesome 4333Font Awesome is fully open source and GPL font friendly.
-https://fontawesome.com, the same font used to design the existing Bitwarden’s iconography.
-We chose green due to its association with safety in Anglo-American contexts (dix2004human, ) and derived it from Bitwarden’s blue (#175DDC) and grey (#7C7C7C), ending up with the shade (#0BDB0B) using Adobe Color [^4] A palette tool referenced by Tidwell et al.
-(tidwell2010designing, ).
-https://color.adobe.com/create.
-The design process followed three phases: (1) Initial brainstorming with security researchers that resulted in 15 icon designs, later narrowed to 7.
-(2) Contextual tests of the icons within the PMs interface led to feedback on aesthetics, size, and positioning.
-(3) An external survey with 20 users with: (a) an initial preference test without explanation, (b) a preference test informed by the icon’s purpose (harley_2016, ), (c) final ratings in the icon’s intended context.
-A detailed description of the icon design process can be found in Appendix A.
-Figure 2. Formal Verification icon
+We designed an icon to communicate formal verification ( 4 in  Figure˜1), and placed it on all features that could be formally verified: (a) Password vault; (b) Primary password input box; (c) Password generator; (d) Clipboard, by its settings.
+We derived the color green used in the UI from Bitwarden’s blue (#175DDC ) and grey (#7C7C7C ), ending up with the shade (#0BDB0B ).
+The design process followed three phases: (a) Initial brainstorming with security researchers that resulted in 15 icon designs, later narrowed to 7 ( 2 in Figure˜1).
+(b) Contextual tests of the icons within the PMs interface.
+(c) A survey with 20 users with two preference tests [22] and ratings in the icon’s intended context.
 Descriptions of formal verification.
-To explain the role of formal verification in the Password Manager, we initially identified all features amenable to formal verification, as well as the potential locations within the Password Manager interface for displaying the verification icon.
-We developed the explanations in three phases: (1) We began with one-on-one discussions with formal verification researchers, gathering explanations, and finally, removing jargon.
-(2) The second iteration consisted of meeting with the team, gathering feedback , and applying it.
+To explain the role of formal verification in the PM, we began by identifying all verifiable features.
+We developed the explanations in three phases: (1) We began with one-on-one discussions with formal verification researchers, gathering explanations, and removing jargon ( 5 in Figure˜1).
+(2) We then met with the team, gathered feedback, and applied it.
 (3) Finally, after another round of feedback, we reached a consensus and finalized the explanations.
-We compared the base Password Manager (i.e., without any extensions or formal verification) and the extended Password Manager (i.e., with interface updates).
-We performed two pilot interviews to refine the protocol and interview script.
+We compared the base PM (i.e., without any extensions or formal verification) and the extended PM (i.e., with interface updates).
 Structure.
-The user tests were divided into four parts.
-First, we provided users with a brief introduction to the study, and after we asked users to fill out the “Pre-Task Questionnaire” about their experience with Password Managers and demographics.
-All participants were asked to perform everyday tasks in a Password Manager (e.g., save a password).
-And finally, we ask them to fill out the “Final Questionnaire” .
+The user tests were divided into the following parts.
+First, we provided users with a brief introduction to the study with a “Pre-Task Questionnaire” about their experience with PMs and demographics.
+We asked participants to perform everyday tasks in a PM (e.g., save a password).
+And finally, we asked them to fill out the “Final Questionnaire”.
 Each questionnaire was used as a base for semi-structured interviews.
-We asked follow-up questions about the participants’ answers.
-Data was collected in two ways: through questionnaires and observation.
-The extended Password Manager did not have formally verified features.
-As such, we included the following disclaimer: “This is a product in development, and some formally verified features are not fully implemented.”.
+We collected data through questionnaires and observation.
 Recruitment and Participant demographics.
-Our sample comprised 15 participants – 10 for the extended interface and 5 for the control condition.
+Our sample comprised 15 participants, 10 for the extended interface and 5 for standard interface.
 Participants were recruited through the authors’ network and received no compensation.
@@
 The most frequent age group was 25-34 (40%).
-Overall, 60% of users were younger than 34, and 40% were at least 35 years old.
+Overall, 60% of users were younger than 34.
 Gender in our sample was evenly divided.
 Each interview took around 50 minutes, and 90% of participants were unfamiliar with formal verification.
-4.2. Results
-To understand if participants knew the icon’s (see Figure 2) meaning, we asked them to explain it.
-Of the 10 participants, 5 mentioned formal verification, and the other five mentioned concepts related to the security of the Password Manager (e.g., “the icon means that the passwords were safe” (P1)).
+We performed two pilot interviews to refine the protocol and interview script ( 6 in Figure˜1).
+4.2 Results
+To understand if participants knew the icon’s meaning (see  4 in Figure˜1), we asked them to explain it.
+Of the 10 participants, 5 mentioned formal verification, and the other five mentioned concepts related to the security of the PM (e.g., “the icon means that the passwords were safe” (P1)).
 Some users correctly identified the formally verified features.
 Another aspect we explore is whether participants understand which features are formally verified.
-Specifically, we found that: (a) 60% identified the generator and the storage as formally verified; (b) 30% stated that the whole Password Manager was formally verified; when asked why one user stated “I saw the icon in several places” (P2); (c) and, one user (10%) could not explain, stating they did not know.
-Users may be more willing to use a formally verified Password Manager.
-When comparing users’ responses regarding trust in a formally verified Password Manager vs one that is not formally verified, users stated trusting more the formally verified Password Manager.
-4.2.1. Limitations and Motivation for Second User Study.
+Specifically, we found that: (a) 6 out of 10 identified the generator and the storage as formally verified; (b) 3 out of 10 stated that the whole PM was formally verified; when asked why one user stated “I saw the icon in several places” (P2); (c) and, one user could not explain, stating they did not know.
+Users may be more willing to use a formally verified PM.
+When comparing users’ responses regarding trust in a formally verified PM vs one that is not formally verified, users stated trusting more the formally verified PM.
+4.2.1 Limitations and Motivation for Second User Study.
 A key limitation is the small sample size, which may not reflect a diverse population.
@@
 To avoid that issue in our main study, we completely refrain from explaining formal verification; instead, we communicate about the guarantees it provides.
-Another limitation is using a specific Password Manager, which might threaten the study’s generality.
-We argue that it did not influence the participants’ responses significantly as the interface does not appear remarkably different from other browser extension Password Managers (e.g., LastPass (lastpass, ) or 1Password (1password, )).
-However, we still decided not to use any specific Password Manager interface in our main study to facilitate generalization.
-After gathering the main insights from this explorative study, we designed and deployed a large-scale quantitative study informed by this formative study.
-5. Main Study
-To address all the research questions mentioned in Section 1 we designed and conducted a 200-participant online survey to understand users’ perceptions of formal verification.
-Table 1. Percentage of participants that agreed or strongly agreed that the factor would impact their willingness to use a PM.
+Another limitation is using a specific PM, which might threaten the study’s generality.
+So, we decided not to use any specific PM interface in our main study.
+After gathering the main insights from this explorative study, we designed and deployed a large-scale quantitative study.
+5 Main Study
+To address all the RQs mentioned in Section˜2, we designed and conducted a 200-participant online survey to understand users’ perceptions of formal verification ( 8 in Figure˜1).
+Table 1: Percentage of participants that agreed or strongly agreed that the factor would impact their willingness to use a PM.
 Results Factors
@@
 87.0% being easy to use for first-time/beginner users
-5.1. Structure
+5.1 Structure
 First Section.
-One of our goals was to understand if formal verification in a Password Manager affects users’ willingness to use it.
-To understand this, we first had to make sure that users knew what a Password Manager was, as previous studies have shown that some users do not use Password Managers because they do not know that these tools exist (pearman2019people, ; stobert2014password, ; alkaldi2016people, ).
-To explain what a Password Manager is and to situate the participants in our survey, we first present them with a vignette scenario where they read a news story about what a Password Manager is and the role of the primary password.
+One of our goals was to test whether formal verification in a PM affects users’ willingness to use it.
+First, we ensured participants knew what a PM was (prior work suggests users may not be familiar with PMs [41, 50].) To do this, we show participants a vignette scenario where they read a news story about what a PM is.
 After viewing this scenario, we show participants a survey question addressing RQ1.
-We ask “Imagine that you were thinking about using a Password Manager.
-What would make you more willing to use the Password Manager?” We then present participants with different factors to which they respond using a Likert (lazar2017research, ) 5-point agreement scale.
+We ask “Imagine that you were thinking about using a PM.
+What would make you more willing to use the PM?” We then present participants with different factors to which they respond using a 5-point Likert scale  [33].
 Among these factors, we include formal verification.
-The reason we added other factors in addition to formal verification was to: (1) Prevent participants from realizing that the focus of the study was formal verification.
-We did this to prevent common biases where the users are inclined to agree with the researcher (merrett2006reflections, ); (2) Understand if users’ preference for formally verified Password Managers differed significantly from other factors.
-These factors were, for example, that the Password Manager provides support materials such as tutorials and is free (see Table 1).
+The reason we added other factors in addition to formal verification was to prevent participants from realizing that the focus of the study was formal verification and prevent common biases where the users are inclined to agree with the researcher.
+We also wanted to understand if users’ preference for formally verified PMs differed from other factors (see Table˜1).
 One of the factors used was a certification by the “Password Manager Security Group”.
 This factor did not represent an actual entity.
-It was added to understand if participants were inclined to agree that any factor was impactful without understanding or because it sounded good.
-Finally, we omit the term ”formal verification” by using analogies when possible.
+It was added to understand if participants were inclined to agree that any factor was impactful, without understanding, or because it sounded good.
+Finally, we omit the term “formal verification” by using analogies when possible.
 The specific analogy we chose is related to something most participants had to deal with in their education — mathematics.
-Instead of saying that the Password Manager is formally verified, we state that it is ”mathematically correct, that is, its features are as trustworthy as a mathematical proof.” We do this to follow best practices (redmiles2017summary, ) and exclude jargon from our survey as most users are unfamiliar with formal verification.
-Additionally, media outlets have previously used mathematical analogies to explain formal verification to broader audiences, such as the BBC (rubens_2015, ) and Quanta Magazine (hartnett_2020, ).
+Instead of saying that the PM is formally verified, we state that it is “mathematically correct, that is, its features are as trustworthy as a mathematical proof.” We do this to follow best practices [46] and exclude jargon from our survey, as most users are unfamiliar with formal verification.
+Media outlets have previously used mathematical analogies to explain formal verification to broader audiences (e.g., BBC [47] and Quanta Magazine [23]).
 Second Section.
-After asking participants if they found that formal verification (among other factors) would impact their willingness to use a Password Manager, we asked them to elaborate.
-We presented participants with their answers and asked, respectively, ”In the previous question, you stated you felt more/less willing to use a Password Manager that is mathematically correct (…).
-Please state your reasons.”.
-We hoped to understand why participants valued (or not) formal verification in a Password Manager and what they associated it with.
-We also hoped to gauge if there was any misconception about what formal verification was for participants.
+After asking participants if they found that formal verification (among other factors), we presented participants with their answers and asked, respectively, “In the previous question, you stated you felt more/less willing to use a PM that is mathematically correct (…).
+Please state your reasons.” We hoped to understand why participants valued (or not) formal verification in a PM and what they associated it with.
 Third Section.
-To answer RQ2 and RQ3, we gather common Password Manager features (e.g., Password Generator and Clipboard clearing).
+To answer RQ2 and RQ3, we gather common PM features (e.g., Password Generator and Clipboard clearing).
 For each of these, we present scenarios representing the impact formal verification can have on each feature.
-For example, for the Password Generator, the policy compliance scenario is: ”Imagine that you are creating a new account on a website (e.g., Twitter, Facebook).
-To increase security, you ask the Password Manager to generate a password with seven characters and at least two numbers.
+For example, for the Password Generator, the policy compliance scenario is: “Imagine that you are creating a new account on a website (e.g., Twitter, Facebook).
+To increase security, you ask the PM to generate a password with seven characters and at least two numbers.
 However, the password generated does not include any numbers.”.
-After each scenario, and using a 5-point Likert (lazar2017research, ) scale, we ask users if that scenario would make them stop using a Password Manager.
+After each scenario, and using a 5-point Likert scale, we ask users if that scenario would make them stop using a PM.
 A feature may have multiple scenarios if formal verification can impact it differently.
 An example is, again, the password generator, where formal verification may help with policy compliance but also with the guarantee that the generator is truly unpredictable (i.e., all passwords have the same probability).
-The scenarios were developed iteratively and followed guidelines for vignette scenarios design (barter1999use, ).
-In total, we designed seven unique scenarios.
-5.2. Recruitment
+We followed design guidelines [33] to iteratively develop seven unique scenarios
+5.2 Recruitment
 We recruited participants through Prolific.
@@
 We recruited 200 participants who were paid 2.65 GBP per submission (with one submission per user and an average time between 8 and 9 minutes).
-We deployed the with LimeSurvey.
-[^6] LimeSurvey is an open-source online survey tool https://www.limesurvey.org/.
-Participant demographics
+Participant demographics.
 About half of our sample identified as male and half as female.
-Our sample skewed younger than the general population as most participants were less than 34 years (87%), and the average age was 26.
-Most participants knew what a Password Manager was before the study (77.5%), but a slight majority of them stated that they did not use one currently (52%).
-The rest of the users were divided between those who said they were currently using a Password Manager (40%), those who had used one in the past but not anymore (3.5%), and those who did not know (3.5%).
-The three Password Managers that participants mentioned the most were Google Chrome Password Manager (41%), Bitwarden (15%), and Apple Keychain (14%).
-Nonetheless, due to a lack of understanding about what a Password Manager is some participants could answer that they had never used a Password Manager when they used a built-in Password Manager (e.g., when they received prompts to save a password in a browser).
-So, to understand the actual number of participants that use a Password Manager, we also asked participants if they saved their passwords in built-in Password Managers.
-As stated before, 52% of participants claimed they had never used a Password Manager before.
-However, of that 52%, 67% admitted to saving passwords when prompted (i.e., using built-in Password Managers).
-These results seem to imply that users use Password Managers without realizing they are using them.
-Reasons for these results could be a lack of technical knowledge, low computer literacy, and a lack of understanding about a Password Manager.
-We do not know if using a Password Manager without knowing impacts users’ perception of these tools.
-In total, 77% of participants used Password Managers to store and manage passwords.
-5.3. Analysis
-To understand if there was a significant difference between different survey questions, we used non-parametrical statistical tests as recommended by Lazar et al.
-(lazar2017research, ).
-We used Friedman’s ANOVA and Wilcoxon Signed-Rank Test with Bonferroni continuity correction (for repeated measures (lazar2017research, )).
-Non-nonparametric tests were used in previous studies on user perceptions of security issues (previous-usenix-nonparametrical, ).
-We coded the open answers from the survey using an inductive coding strategy.
-Two coders created the codebook iteratively using inductive and hierarchical coding (lazar2017research, ).
-In total, two coders coded 200 answers.
-We used Cohen’s Kappa to ensure inter-rater reliability and finished the coding process with a Cohen’s Kappa of 0.81, which denotes a very high level of agreement.
-An interpretation of Cohen’s Kappa is that a value above 0.60 indicates “satisfactory reliability” and one above 0.80 ”near-perfect agreement”  (lazar2017research, ).
-All answers were coded individually.
-However, regular meetings happened to discuss ambiguities and possible changes to the codebook.
-The final codebook consisted of 13 codes (see Appendix H), and each participant’s answer could have multiple codes.
-We used a hierarchical coding scheme for some of the codes, for example, the code “Extra Security” had as sub-codes: “Extra security: password generation”, for when participants mentioned that the Password Manager could create strong passwords; “Extra security: secure storage”, for when the participant stated that their passwords were safer inside the Password Managers’ storage; and ”Extra security: Protection from unwanted access”, for when they mentioned that the Password Manager prevents third-parties from accessing their passwords.
-We also included more general codes for when participants were not specific in their answers or showed a lack of understanding about what was being asked (”Answer general to all Password Managers and not specific to formally verified Password Managers” and ”Just seems better”) and one code for answers that were unusable to us (”Answer not relevant”).
-Two answers were coded as unusable, leaving this study with 198 usable participant answers for this question.
-Figure 3. Frequency of codes representing the reasons participants gave for ”agreeing” or ”disagreeing” that they felt willing to use a mathematically correct Password Manager.
-5.4. Results
+Our sample skewed younger than the general population, as most participants were less than 34 years (87%), and the average age was 26.
+Most participants knew what a PM was before the study (77.5%), but a slight majority of them stated that they did not use one currently (52%).
+The rest of the users were divided between those who said they were currently using a PM (40%), those who had used one in the past but not anymore (3.5%), and those who did not know (3.5%).
+The three PMs that participants mentioned the most were Google Chrome PM (41%), Bitwarden (15%), and Apple Keychain (14%).
+Nonetheless, due to a lack of understanding about what a PM is, some participants answered that they had never used a PM when they used a built-in PM (e.g., when they received prompts to save a password in a browser).
+We saw that, of the 52% of participants who had never used a PM, 67% of them saved passwords when prompted (e.g., in built-in PMs).
+These results seem to imply that users use PMs without realizing they are using them, maybe due to a lack of technical knowledge or low computer literacy.
+In total, 77% of participants used PMs to store and manage passwords.
+5.3 Analysis
+To understand if there was a significant difference between different survey questions, we used Friedman’s test and Wilcoxon Signed-Rank Test with Bonferroni continuity correction [33].
+We coded the open answers from the survey using an inductive coding strategy ( 12 in Figure˜1).
+Two coders created the codebook iteratively using inductive and hierarchical coding [33].
+We coded 200 answers, achieving a Cohen’s Kappa of 0.81, indicating very high inter-rater reliability [33].
+All answers were coded individually, and regular meetings happened to discuss ambiguities and updates to the codebook.
+The final codebook consisted of 13 codes and each participant’s answer could have multiple codes.
+We used a hierarchical coding scheme for some of the codes, for example, the code “Extra Security” had as sub-codes: “Extra security: password generation”, for when participants mentioned that the PM could create strong passwords and “Extra security: secure storage”, for when the participants stated that their passwords were safer inside the PMs’ storage.
+We also included more general codes for when participants were not specific in their answers or showed a lack of understanding about what was being asked (“Answer general to all PMs and not specific to formally verified PMs” and “Just seems better”).
+Figure 2: Frequency of codes representing the reasons participants gave for “agreeing” or “disagreeing” that they felt willing to use a mathematically correct PM.
+5.4 Results
 In this section, we address each RQ and its respective study insights.
-5.4.1. Formal Verification and Willingness to use a Password Manager (RQ1)
-In the survey’s first question, we asked participants whether formal verification impacted their willingness to use a Password Manager.
-To prevent biases, we compared seven positive factors (see Table 1).
-The factors are all things that could be perceived as desirable to participants, so it is no surprise that most participants ”agreed” or ”strongly agreed” would be more likely to use a Password Manager with any of those factors.
+5.4.1 Formal Verification and Willingness to use a PM (RQ1)
+In the survey’s first question, we asked participants whether formal verification impacted their willingness to use a PM.
+To reduce biases, we compared seven positive factors (see Table˜1).
+The factors are all things that could be perceived as desirable to participants, so it is no surprise that most participants “agreed” or “strongly agreed” would be more likely to use a PM with any of those factors.
 What we want to understand is if their opinion is the same across all factors or if there are some factors that users value more.
-A Friedman test showed that there was a statistically significant difference in participants’ answers between the different Password Manager’s features with a $\chi^{2}(2)=51.42$ and a $p<0.05$.
+A Friedman test showed that there was a statistically significant difference in participants’ answers between the different PM’s features with a $\chi^{2}(2)=51.42$ and a $p<0.05$.
 However, the following pairwise comparison of users’ answers using the Wilcoxon rank sum test did not find a significant difference between most factors ($p>0.05$).
-Comparing all the factors, we found that more than 86% of participants considered the use of formal verification (as described in Section 5.1) and being easy to use as important (they ”agreed” or ”strongly agreed” that these factors would make them more willing to use a Password Manager).
-The factors that users considered less critical (while still being essential for most) were being inexpensive (i.e., it costs money even if it is not expensive) and having support features (e.g., tutorials and help pages) with 31% and 27%, respectively, not thinking they would be more willing to use a Password Manager with them.
-This information can be found in Table 1.
-Our results suggest that users are more willing to use a formally verified Password Manager than a non-verified one.
+Comparing all the factors, we found that more than 86% of participants considered the use of formal verification (as described in Section˜5.1) and being easy to use as important (they “agreed” or “strongly agreed” that these factors would make them more willing to use a PM).
+The factors that users considered less critical (while still being essential for most) were being inexpensive (i.e., it costs money even if it is not expensive) and having support features (e.g., tutorials and help pages) with 31% and 27%, respectively, not thinking they would be more willing to use a PM with them.
+This information can be found in Table˜1.
+Our results suggest that users are more willing to use a formally verified PM than a non-verified one.
 With this information in mind, we now aim to understand why participants answered our survey the way they did.
-Understanding User’s Reasons.
-This section analyzes the second survey section and aims to understand users’ reasons for varying formal verification.
-We separated the coding into the three questions that could be asked to participants — participants that valued formal verification, had no opinion, and found it undesirable (green, grey, and red in Figure 3).
-Additionally, depending on the size of the participant’s answer, one or more codes could be applied.
-Most participants who valued formal verification in their Password Managers mentioned the use of mathematics as a reason why (67 participants (35%)).
-Participants seem to associate mathematics with a “logical” and certain behavior.
-One even stated that they preferred to use a formally verified Password Manager “because mathematics doesn’t lie” (P89).
+Understanding Users’ Reasons.
+We separated answers into three categories: participants who valued formal verification, had no opinion, and found it undesirable (see Figure˜2).
+Most participants who valued formal verification in their PMs cited the use of mathematics as a reason (67 participants (35%)).
+Participants seem to associate mathematics with a “logical” and precise behavior.
+One even stated that they preferred to use a formally verified PM “because mathematics doesn’t lie” (P89).
 Another stated, “I like to use such products or services that have scientific proof of their effectiveness” (P147).
-“Security” codes were also frequent (see Section 5.3).
-Some participants mentioned they believe formal verification added security (e.g., ”I think it would be more secure.” (P91)), and others gave more detailed explanations (e.g., “The main reason to start using a password manager is (…) the passwords generated are the strongest (…)” (P5)).
-Over 50% of the participants mentioned security as a reason why they valued formally verified Password Managers, with some mentioning more than one specific subcode of security in their answers (e.g., one participant stated that they preferred a formally verified Password Manager because ”…it will be more difficult to get hacked since it will create more secure passwords” (P162)).
-Overall, participants mostly agreed that formal verification was something to be desired in a Password Manager.
-They provided several reasons, but the more frequent ones were related to security and the use of mathematics.
-On the other hand, participants who did not find formal verification relevant had a general distrust of Password Managers, found that formal verification was not essential for them, or simply did not understand the concept.
-For example, one of the participants directly mentioned, ”I just don’t believe in that kind of app.
-” (P22).
-Lack of understanding was also a frequent code present in the answers of participants who had a neutral response to this survey question (see the bars in yellow in Figure 3).
-Lack of understanding of technical concepts has been the cause of users’ trust issues with other security software in the past (e.g., in the domain of cryptocurrencies, lack of knowledge on the subject may cause users to question the security of these systems (presthus2017motivations, )).
+“Security” codes were also frequent (see Section˜5.3).
+Some participants mentioned they believe formal verification added security (e.g., “I think it would be more secure.” (P91)), and others gave more detailed explanations (e.g., “The main reason to start using a PM is (…) the passwords generated are the strongest (…)” (P5)).
+Over 50% of the participants mentioned security as a reason why they valued formally verified PMs, with some mentioning more than one specific subcode of security in their answers (e.g., one participant stated that they preferred a formally verified PM because “… it will be more difficult to get hacked since it will create more secure passwords” (P162)).
+Overall, participants mostly agreed that formal verification was something to be desired in a PM.
+The most frequent reasons for this were related to security.
+On the other hand, participants who did not find formal verification relevant had a general distrust of PMs or found that formal verification was not essential for them.
+For example, one of the participants directly mentioned, “I just don’t believe in that kind of app” (P22).
+Lack of understanding was also a frequent code present in the answers of participants who had a neutral response to this survey question (see the bars in grey in Figure˜2).
+Lack of understanding of technical concepts has been the cause of users’ trust issues with other security software in the past [44].
 Nonetheless, most participants who agreed that formal verification was important to them stated that security had something to do with it.
 RQ1.
-How does formal verification impact users’ willingness to use Password Managers?
-Users seem to be more willing to use a formally verified Password Manager than a non-formally verified Password Manager; Our results also suggest that formal verification has positively affected some users’ trust; The main reasons users value formal verification are the use of mathematics and the extra security formal methods provide.
-Table 2. P-values of pairwise comparisons of users’ answers in the scenarios question using the Wilcoxon rank sum test and Bonferroni continuity correction. See Table 4 for the scenario descriptions.
+How does formal verification impact users’ willingness to use PMs?
+Users seem to be more willing to use a formally verified PM than a non-formally verified PM; Formal verification seems to have positively affected some users’ trust; Users primarily value formal verification for its reliance on mathematics and the additional security it provides.
+5.4.2 Formally Verified Features (RQ2)
+After learning that formal verification seems to impact users’ interest in using PMs, we are interested in learning their priorities, thus addressing RQ2.
+We realize that users may not have the expertise to judge what is essential and what should be verified in a PM.
+However, our goal is to understand what features, if verified, would be more impactful for users.
+We want to understand how to meet users’ wants and maximize the impact of formal verification.
+A Friedman’s test showed that there was a statistically significant difference in participants’ answers between the different scenarios with a $\chi^{2}(2)=331.08$ and a $p<0.05$.
+The results of a pairwise comparison of users’ answers using the Wilcoxon rank sum test and Bonferroni continuity correction can be seen in Table˜2.
+Our results suggest that most of the differences between scenarios are statistically relevant ($p<0.05$).
+Table 2: P-values of pairwise comparisons of users’ answers in the scenarios question using the Wilcoxon rank sum test and Bonferroni continuity correction.
 Scenario 1 Scenario 2 Scenario 3 Scenario 4 Scenario 5 Scenario 6 Scenario 7
@@
 Scenario 8 1.1e-14 0.00138 2e-16 1.8e-11 0.00024 0.00019 0.08562
-5.4.2. Formally Verified Features (RQ2)
-After learning that formal verification seems to impact users’ interest in using Password Managers, we are interested in learning their priorities, thus addressing RQ2.
-We realize that users may not have the expertise to judge what is essential and what should be verified in a Password Manager.
-However, our goal is to understand what features, if verified, would be more impactful for users.
-We want to understand how to meet users’ wants and maximize the impact of formal verification.
-A Friedman’s test showed that there was a statistically significant difference in participants’ answers between the different scenarios with a $\chi^{2}(2)=331.08$ and a $p<0.05$.
-The results of a pairwise comparison of users’ answers using the Wilcoxon rank sum test and Bonferroni continuity correction can be seen in Table 2.
-Our results suggest that most of the differences between scenarios are statistically relevant ($p<0.05$).
-Table 3. Percentage of participants that agreed or strongly agreed the scenario would make them stop using a PM. S# corresponds to Scenario # (see Table 4 in the Appendix for scenario descriptions).
+Table 3: Percentage of participants that agreed or strongly agreed the scenario would make them stop using a PM. S# corresponds to Scenario #
 Results Scenarios
@@
 Overall, participants seem to consider some features more critical than others.
-Our results suggest that the secure vault and its encryption are the most important – 96% of participants stated they agreed or strongly agreed that they would stop using a Password Manager if the respective scenario (S3) occurred.
-In order the most important features to verify seem to be the secure vault (S3), password generator (S1 and S2), login in the Password Manager (i.e., primary password security in S4), clipboard (S6), synchronization across devices (S7 and S8) – see Table 3.
+Our results suggest that the secure vault and its encryption are the most important — 96% of participants stated they agreed or strongly agreed that they would stop using a PM if the respective scenario (S3) occurred.
+The most important features to verify seem to be the secure vault (S3), password generator (S1), login in the PM (i.e., primary password security in S4), clipboard (S6), synchronization across devices (S7 and S8) (see Table˜3).
 RQ2.
-What features would users like to see formally verified in a Password Manager?
-The password vault’s security seems to be the feature that should be prioritized in a future implementation of a formally verified Password Manager; Surprisingly, verifying that the Password Manager’s passwords will not be lost does not seem to be important for users, as scenarios, where there was no third-party access to the password, were not as impactful; Ideally, all the features mentioned in this section should be formally verified, as over 50% of participants found their scenarios impactful.
-5.4.3. Formal Verification Guarantees (RQ3)
-The scenarios’ question also provides insights on whether participants value the guarantees of formal verification in Password Managers, thus answering RQ3.
+What features would users like to see formally verified in a PM?
+The password vault’s security seems to be the feature that should be prioritized in a future implementation of a formally verified PM; Surprisingly, verifying that the PM’s passwords will not be lost does not seem to be important for users, as scenarios where there was no third-party access to the password were not as impactful; Ideally, all the features mentioned in this section should be formally verified, as over 50% of participants found their scenarios impactful.
+5.4.3 Formal Verification Guarantees (RQ3)
+The scenario’s question also provides insights into whether participants value the guarantees of formal verification in PMs.
 As stated before, most scenarios were impactful for users, so in general, our results suggest that participants value the guarantees that formal verification provides.
-Additionally, we can divide our scenarios into the following categories:
-  • Scenarios where a third-party learns all the passwords: scenarios ”Vault Exposure” (S3) and ”Primary Password Exposure” (S4);
-  • Scenarios where a third party, over time, can learn a large number of passwords: scenarios ”Unpredictability” (S1), ”Autofill” (S5), and ”Clipboard Clearing” (S6);
-  • Scenarios where the user loses access to all the passwords stored in the Password Manager’s vault: scenarios ”Synchronization” (S7) and ”Deleting your vault” (S8);
-  • A scenario where the generator does not behave as intended and the passwords generated are not compliant with the users’ inputs: scenario ”Policy Compliance” (S2).
-Two of the three most impactful scenarios described a situation where a third party could learn all the users’ passwords (S3 and S4).
-In these scenarios, 86.5% and 92.5% of participants stated they ”agreed” or ”strongly agreed” that they would stop using a Password Manager if it happened.
-These scenarios were related to the absence of formal verification where its guarantees would prevent the exposure of passwords to a third party.
+We can divide our scenarios into the following categories:
+  • Scenarios where a third-party learns all the passwords: “Vault Exposure” (S3) and “Primary Password Exposure” (S4);
+  • Scenarios where a third party, over time, can learn manypasswords: “Unpredictability” (S1), “Autofill” (S5), and “Clipboard Clearing” (S6);
+  • Scenarios where the user loses access to all the passwords stored in the PM’s vault: “Synchronization” (S7) and “Deleting your vault” (S8);
+  • A scenario where the generator does not behave as intended and the passwords are not compliant with the specification: “Policy Compliance” (S2).
+Two of the three most impactful scenarios allowed a third party to learn all user passwords (S3, S4).
+In these scenarios, 86.5% and 92.5% of participants stated they “agreed” or “strongly agreed” that they would stop using a PM if it happened.
+These scenarios involved the lack of formal verification, which would otherwise prevent password exposure to third parties.
 However, the second, fourth, and fifth most impactful scenarios (S1, S5, and S6) described a situation where a third party, over time, could learn several users’ passwords.
-In these, over 79% of participants ”agreed” or ”strongly agreed” that they would stop using a Password Manager if they happened.
-The other type of scenario, where the user loses access to all the passwords they have stored in the Password Manager’s vault (S7 and S8), was not as impactful for participants as the ones mentioned before.
-Finally, the last type of scenario described a situation where the Password Manager did not work as intended (S2), but its malfunction did not lead to a loss of passwords.
+In these, over 79% of participants “agreed” or “strongly agreed” that they would stop using a PM if they happened.
+The other type of scenario, where the user loses access to all the passwords they have stored in the PM’s vault (S7 and S8), was not as impactful for participants as the ones mentioned before.
+Finally, the last type of scenario described a situation where the PM did not work as intended (S2), but its malfunction did not lead to a loss of passwords.
 It just inconvenienced users.
 This was the scenario that fewer participants found impactful.
-Interestingly, the least important feature in our hierarchy—the unpredictability of the password generator (S1)—is a feature that previous work has made efforts to formally verify (grilo22pwdgen, ).
-Our results thus suggest that users value the guarantees that formal verification provides.
-Moreover, users value the guarantees more depending on the impact that said guarantees have.
-Guaranteeing that passwords are not leaked to third parties (e.g., by ensuring the cryptographic algorithms are well implemented) seems to be very important for participants.
-RQ3: Do users value the guarantees that formal verification can provide in Password Managers?
-Our results suggest that users value the guarantees that formal verification can provide in Password Managers; Our data also seems to indicate that some guarantees are more important than others.
+Interestingly, the least important feature in our hierarchy, the unpredictability of the password generator (S1), is a feature that previous work has made efforts to formally verify [20].
+RQ3: Do users value the guarantees that formal verification can provide in PMs?
+Our results suggest that users value the guarantees that formal verification can provide in PMs; Our data also seems to indicate that some guarantees are more important than others.
 For example, guaranteeing that the passwords are not leaked to third parties seems more relevant for participants than guaranteeing that they are synced in the cloud.
-6. Discussion
-In this section, we situate our results within the broader context of formal verification application.
-We discuss the main insights and suggest recommendations for future verification efforts and research.
-Our main insights include:
+6 Discussion
+In this section, discuss the main insights and suggest recommendations.
 User Awareness and Perception.
 An insight we got was that there is an overall low level of awareness among users about what formal verification is.
-Despite this, users react positively to formal verification in Password Managers.
-There appears to be a gap between the technical understanding of formal verification and the perceived value it adds to Password Managers.
+Despite this, users react positively to formal verification in PMs.
+There appears to be a gap between the technical understanding of formal verification and the perceived value it adds to PMs.
 Some of our participants liked formal verification due to an inherent trust in the mathematical and technical rigor it implies, suggesting that even a superficial understanding or awareness of formal verification may influence user preferences.
 User features prioritization.
-A significant insight from this study is that users perceive some Password Manager features as more important.
+A significant insight from this study is that users perceive some PM features as more important.
 The security of the password vault and the reliability of password generation were highlighted as particularly critical features for users.
 This insight underscores the need for developers to prioritize formal verification efforts on features that most directly impact user trust and perceived security.
-Understanding user priorities can guide the allocation of development resources towards the aspects of Password Managers that users value more and enhance the effectiveness and attractiveness of formally verified Password Managers.
-6.1. Recommendations
+Understanding user priorities can guide the allocation of development resources towards the aspects of PMs that users value more and enhance the effectiveness and attractiveness of formally verified PMs.
+6.1 Recommendations
 One of our goals was to identify how to increase the impact of formal verification in software.
@@
 We suggest that these efforts should be focused on the specific consequences of formal verification in software instead of theoretical concepts.
-Avoiding technical language may facilitate the application of best practices such as using metaphors and avoiding jargon (redmiles2017summary, ).
+Avoiding technical language may facilitate the application of best practices such as using metaphors and avoiding jargon [46].
 Increasing this transparency may enable users to understand the advantages of formally verified software and, thus, increase its adoption.
-If formally verified software is valued by users, it may also increase its investment value for businesses (for example, for Password Managers such as Bitwarden (bitwarden, )).
+If formally verified software is valued by users, it may also increase its investment value for businesses (for example, for PMs such as Bitwarden).
 2. Don’t overstate/mislead users when explaining formally verified software.
-As mentioned before, some users associated formal verification with security (see Section 4.2).
-While this sometimes can be true, security is not necessarily related to formal verification.
+As mentioned before, some users associated formal verification with security (see Section˜4.2).
+While this can sometimes be true, security is not necessarily related to formal verification.
 As such, we suggest that future efforts to communicate about formal verification should take care to prevent overstating the impact it has.
-Misleading users can be counterproductive and lead to distrust.
+To build trust, communication efforts should carefully distinguish the benefits of verification from broader security properties.
 3. Make an effort to understand the users’ priorities in formally verified software.
@@
 Moreover, since formally verifying software is time-consuming and expensive, understanding users’ priorities may help focus verification efforts on the features that are more impactful.
-6.2. Limitations
+6.2 Limitations
 User studies such as these may suffer from bias.
 Bias can arise from the questions or even the questionnaires.
-To mitigate our limitations, we performed cognitive interviews [^7] Cognitive interviews involve asking respondents to think aloud as they complete a survey and asking them questions about each survey item (redmiles2017summary, ).
+To mitigate our limitations, we performed cognitive interviews [^3] Cognitive interviews involve asking respondents to think aloud as they complete a survey and asking them questions about each survey item [46].
 to validate the study until no more errors or typos were detected (we did two cognitive interviews for each study).
-We also followed best practices by offering “don’t know (DK)” or “prefer not to answer” responses (redmiles2017summary, ).
-Moreover, we asked users to explain their understanding of the topic to mitigate the Dunning-Kruger Effect (e.g., when they stated that they understand what formal verification is, we asked them to explain what it is) and included attention check questions.
-For example, we include one of these questions among the factors that may impact users’ willingness to use a Password Manager (see Section 5.1) and another among the scenarios in the third section of the survey (see Section 5.1).
+We also followed best practices by offering “don’t know (DK)” or “prefer not to answer” responses [46].
+Moreover, we asked users to explain their understanding of the topic to mitigate the Dunning-Kruger Effect (e.g., when they stated that they understood what formal verification is, we asked them to explain what it is) and included attention check questions.
+For example, we include one of these questions among the factors that may impact users’ willingness to use a PM (see Section˜5.1) and another among the scenarios in the third section of the survey (see Section˜5.1).
 We also removed as much jargon as possible.
-The jargon includes terms like “memory” and “encrypted” but also “formal verification” itself (redmiles2017summary, ) and tried to mitigate the Hawthorne effect [^8] Under the Hawthorne effect users are inclined to agree with researchers (merrett2006reflections, ).
-by hiding that the study is about formal verification.
-To prevent bias we randomized the order in which the factors that may impact users’ willingness to use a Password Manager are shown (see Section 5.1), and the order of the scenarios in the third survey section (see Section 5.1).
-6.3. Future Work
-Exploring formal verification’s impact on users’ willingness to use Password Managers opens up several avenues for further research.
+Jargon includes terms like “memory”, “encrypted” and “formal verification” itself [46] and tried to mitigate the Hawthorne effect.
+To prevent bias, we randomized the order in which the factors that may impact users’ willingness to use a PM are shown (see Section˜5.1), and the order of the scenarios in the third survey section (see Section˜5.1).
+6.3 Future Work
+Exploring formal verification’s impact on users’ willingness to use PMs opens up several avenues for further research.
 Exploration in Other Domains.
-Our study within the context of Password Managers suggests a positive user reception towards formal verification.
+Our study within the context of PMs suggests a positive user reception towards formal verification.
 Extending this research to other domains where formal verification is applied could offer an insightful research opportunity.
 For example, investigating user perceptions in the context of autonomous vehicles, medical devices, and blockchain technologies — all of which rely on the integrity and security assurances that formal verification provides — could reveal domain-specific user attitudes and expectations.
-These studies could help understand whether the positive inclination towards formal verification observed in Password Manager users is universal or if adjustments in communication and implementation strategies are needed based on the application domain.
+These studies could help understand whether the positive inclination towards formal verification observed in PM users is universal or if adjustments in communication and implementation strategies are needed based on the application domain.
 Development of New Ways to Communicate about Formal Verification.
 One significant finding from our study is the general lack of user awareness about formal verification.
-This gap presents a research opportunity to develop and test new communication methods about formal verification to the public.
-Inspired by initiatives like the IoT security labels(emami2020ask, ), a “Formal Verification Label” could be designed to provide at-a-glance information about the verification status of a product.
+This gap presents a research opportunity to develop and test new communication methods about formal verification for the public.
+Inspired by initiatives like the IoT security labels [13], a “Formal Verification Label” could be designed to provide at-a-glance information about the verification status of a product.
 Such a label could include simplified symbols or ratings that indicate the extent and areas of formal verification applied, making it easier for users to understand the claims of formally verified software products.
-7. Conclusion
-Formal software verification can be expensive and time-intensive, but our work suggests that it may positively impact users.This insight can be a powerful motivation for future formal verification efforts.
+7 Conclusion
+Formal software verification can be expensive and time-intensive, but our work suggests that it may positively impact users.
 In our study, we propose several directions for future research while providing concrete insights into which features should be prioritized by practitioners, such as enhancing the security of the password vault and ensuring the reliability of password generation.
-By focusing on these priorities, practitioners can allocate resources more efficiently, ensuring their efforts directly address the aspects of Password Managers that users value.Moreover, our work has shed light on a previously mostly unexplored area of research— combining formal verification with usable security.
-We argue that many paths for future work could be explored: virtually any domain where formal methods could be applied has to gain from studying user perception on the subject.
-8. Acknowledgments
-This work was funded by Fundação para a Ciência e a Tecnologia (FCT) under grant PRT/BD/153739/2021, and projects UIDB/50021/2020 (DOI: 10.54499/UIDB/50021/2020), LA/P/0063/2020 (DOI: 10.54499/LA
-/P/0063/2020), and the PassCert project, a CMU Portugal Exploratory Project with reference CMU/TIC/0006/2019.
-Appendix A Design of the Formal Verification Icon
-Design of the Formal Verification Icon
-The icon for formal verification should match the interface of Bitwarden, and because of that, we used the same font to design it (Font Awesome 4999Font Awesome is fully open source and GPL font friendly.
-https://fontawesome.com).
-Color is also very important: one of its roles is to contribute to the aesthetic value of the interface, but it can also be used to transmit information and influence the user emotionally (shneiderman2016designing, ).
-Although it is difficult to assume any universal interpretation of color, green is associated with safety in the Anglo-American cultural color convention (dix2004human, ).
-So, to differentiate from Bitwarden’s interface and draw attention to the formal verification logo, we used a new color, specifically green.
-Bitwarden’s icons use two main colors, Bitwarden’s blue (#175DDC) and grey (#7C7C7C).
-Using a quadratic schema from Bitwarden’s blue and with the help of the Adobe Color [^10] A palette tool referenced by Tidwell et al.
-(tidwell2010designing, ).
-https://color.adobe.com/create, we derived the shade of green used in the icon (#0BDB0B).
-It is also important to consider the needs of color-blind individuals.
-They may confuse some shades of orange or red with green or not see a red dot on a black background (shneiderman2016designing, ).
-Currently, in the Bitwarden interface, we do not have any colors that could be mistaken by color-blind users.
-Although we use green for the Bitwarden symbol, it remains legible (it is not a green text in a red background, for example).
-The development of the icon went through three phases: (1) Together with a small group of colleagues developing research in security and formal methods, we began by brainstorming icon designs and ideas.
-After reaching 15 different variations, we had a group meeting to gather feedback.
-At this stage, we shortlisted 7 icons.
-(2) We then placed the 7 icon variations in their context (in the Password Managers interface), met again with the group, and asked for feedback.
-The team gave specific feedback about icon size, subjective aesthetic value, icon position in the tool, places where the icon should be, and lastly, about what icons represented more accurately formal verification.
-(3) To gather more unbiased feedback, we asked 20 users outside the team to give their opinions.
-Because the icon is intended to mean that a certain feature is formally verified, it is very important to the underlying message we want to transmit.
-As such, we performed user studies to determine the best variation of the icon.
-The participants were recruited through the authors’ personal network and did not receive compensation.
-The form where we asked for feedback about the icons was divided into three parts: (a) First, we asked users to choose the icons they liked more without knowing what they were meant to represent; (b) Secondly, we did a preference test, where we explained what the icon is trying to convey and asked users to again rate the icons according to the ones they prefer (harley_2016, ); (c) And lastly, a preference test, where we showed the icons in the context they would be in the final interface and asked users to rate them.
-This survey can be viewed in Appendix B, and the icon chosen can be seen in Figure 2.
-After an internal discussion about whether the icon could be confused with the HTTPS padlock, we argue that a padlock is not uniquely linked with the HTTPS security icon (von2022builds, ).
-Several icons were developed, and the final design was informed by several feedback meetings with the research team and a thorough user test, as described.
-Appendix B Icon Survey
-The survey used in the icons’ user study is divided in three main parts: Attractiveness Test (§B.2), Preference Test with no context (§B.3), and Preference Test with context (§B.4).
-The survey starts with the participants going through the informed consent form (§B.1).
-B.1. Consent
-The goal of this survey is to collect feedback on a set of icons that are developed to improve the user interface of password managers.
-The icons aim to convey that a certain component is formally verified (this is explained below).
-This project is part of the PROJECT_NAME research project.
-All the data collected is anonymous and will be used solely by the researchers of PROJECT_NAME.
-The data may be used to present insights at conferences, academic events, or similar events and for scientific publications.
-Your participation is voluntary, and you may always quit at any time, without any kind of penalization.
-By selecting “Yes” below, you are consenting for your data to be processed, stored, and used as described above.
-You also are confirming that you have read this consent form.
-The session is predicted to take about 4 minutes.
-B.2. Part 1 – Attractiveness Test
-Please order these icons from the one you find more aesthetically pleasing to the one you find least aesthetically pleasing (for example: E-G-F-H-I-K-J)
-Figure 4. Icon options
-B.3. Part 2 – Preference Test without context
-Formal verification is a process in which developers prove (mathematically) that a certain part of a program behaves as intended.
-To say that a feature is formally verified means that it is guaranteed to behave a certain way.
-A concrete example is guaranteeing that the length of generated passwords satisfies the minimum requirements.
-Order these icons from the one you find that represents best the concept ”Formally Verified Feature” to the one that represents it worst (for example: E-G-F-H-I-K-J)
-Figure 5. Icon options
-B.4. Part 3 – Preference Test with context
-Formal verification is a process in which developers prove (mathematically) that a certain part of a program behaves as intended. The icons bellow are meant to represent this concept.
-1.
-Considering the previous information, please order these images from the ones you like more and that represent the concept better to the worst (for example: E-G-F-H-I-K-J)
-Figure 6. Icon options in the PM’s interface
-2.
-Considering the previous information, please order these images from the ones you like more and that represent the concept better to the worst (for example: E-G-F-H-I-K-J)
-Figure 7. Icon options in the PM’s interface
-3.
-Considering the previous information, please order these images from the ones you like more and that represent the concept better to the worst (for example: E-G-F-H-I-K-J)
-Figure 8. Icon options in the PM’s interface
-4.
-Thank you for filling this survey. Please use this space to give any feedback you consider relevant
-Appendix C First User Study: Interview Protocol
-This appendix shows the interview protocol followed for the first user study. Each of the nine steps includes the script followed.
-1.
-Introduce yourself. Give them the consent form to sign, on their worksheet. If it is not signed, do not proceed.
-My name is NAME, and I will be giving you instructions on what to do and will answer your questions.
-In this questionnaire the participant should read and accept the following consent form:
-“The goal of this survey is to information about you, your experience with Password Managers and specifically with PROJECT_NAME’s Password Manager.
-This project is part of the PROJECT_NAME research project, a project that is building an open-source, proof-of-concept password manager that through the use of formal verification, is guaranteed to satisfy properties on data storage and password generation.
-All the data collected is anonymous and will be used solely by the researchers of PROJECT_NAME.
-The data may be used to present insights at conferences, academic events, or similar events and for scientific publications.
-Your participation is voluntary, and you may always quit at any time, without any kind of penalty.
-By selecting ”Yes” below, you are consenting for your data to be processed, stored, and used as described above.
-You also are confirming that you have read this consent form.”
-2.
-Describe the purpose of the study.
-The goal of this project is to evaluate the interface of PROJECT_NAME’s Password Manager.
-A password manager is a program used to store and manage passwords.
-This is a product in development and some formally verified features are not fully implemented.
-We’re looking for places where the product may be difficult to use.
-If you have trouble with some of the tasks, it’s the product’s fault, not yours.
-Don’t feel bad; that’s exactly what we’re looking for.
-If we can locate trouble spots, then we can go back and improve the product.
-Remember we’re testing the product, not you.
-If you have any feedback we ask you to be as honest as possible.
-3.
-Tell the participant that it’s OK to quit at any time.
-Although I don’t know of any reason for this to happen, if you should become uncomfortable or find this test objectionable in any way, you are free to quit at any time.
-4.
-Explain how to think aloud
-We have found that we get a great deal of information from these informal tests if we ask people to think aloud as they work through the exercises.
-It may be a bit awkward at first, but it’s really very easy once you get used to it.
-All you have to do is speak your thoughts as you work.
-If you forget to think aloud, I’ll remind you to keep talking.
-As you go through these tasks, we ask that you think aloud.
-You can say anything that comes to mind.
-This helps us better understand how users like yourself experience the product.
-If you forget to think aloud, I’ll remind you to keep talking.
-5.
-Explain that you will not provide help.
-As you’re working through the tasks, I won’t be able to provide help or answer questions.
-This is because we want to create the most realistic situation possible.
-Even though I won’t be able to answer your questions, please ask them anyway.
-It’s very important that I capture all your questions and comments.
-When you’ve finished all the exercises, I’ll answer any questions you still have.
-6.
-Time and anonymity.
-The results of this evaluation will be summarized and used in the context of the PROJECT_NAME research project. If the results of this experiment are published your identity will be anonymized.
-The expected duration of this session is 50 minutes.
-Do you have any questions at this point?
-7.
-Give the pre-study questionnaire.
-Remember that a Password Manager is a software used to store and manage passwords.
-Before we begin, please fill out a pre-study questionnaire.
-This will provide us useful information about your demographics.
-Before we start the tasks, I would like you to explore the product. Feel free to click any buttons and take your time.
-8.
-Go through the tasks in the interface. Ask participants to fill the Final Questionnaire.
-Now I would like to ask you to fill the Final Questionnaire.
-9.
-Further feedback. Discuss any interesting behaviors you would like the participant to explain.
-Do you have further feedback or suggestions regarding the product?
-Thank you for your participation.
-If you have any questions or wish to learn more about this research project do not hesitate to contact us.
-Appendix D Pre Task Questionnaire
-The pre-task questionnaire was used as the basis for a semi-structured interview where participants were encouraged to think-aloud about the questions.
-Most Likert Scale survey answers were accompanied by the participants reasons for their answers.
-D.1. Part 1 – Password Managers
-Q1.
-Did you know what a password manager (PM) was before this study?
-If you answered “yes” to the previous question please explain what a Password Manager is.
-Q2.
-Do you use a password manager?
-Choose one of the following answers: Yes, I currently use a PM; Yes, I currently use a PM; No, I’ve never used one; I’m not sure.
-Q3.
-How do you manage/remember passwords?
-D.2. Part 2 - Exclusively for participants that have never used a PM
-Q1.
-Before this study have you ever been interested in using a PM? Choose one of the following answers: Yes; No, and I don’t know what they are; No, but I know what Password Managers are.
-Q2.
-How much do you agree with the following statement:
-I want to use a Password Manager in the future.
-Choose one of the following answers: Strongly disagree, Disagree, Neither agree or disagree, Agree or Strongly Agree.
-D.3. Part 3 - Exclusively for participants that are currently using a PM
-Q1.
-What Password Manager do you use?
-Q2.
-How often have you used your Password Manager (i.e. using the credentials in a Password Manager to log in websites)?
-Q3.
-In which context do you use a Password Manager (e.g. Work, Personal Life, etc)?
-Q4.
-What types of account do you save in the PM?
-Choose all that apply: Social Media, Homebanking, Credit Cards, Online shopping, Mail, News/Entertainment, Other Options (open answer).
-D.4. Part 4 - Exclusively for participants that have stopped using a PM
-Q1.
-What Password Manager did you use?
-Q2.
-Why did you stop using it?
-Q3.
-How much do you agree with the following statements:
-I want to use a Password Manager in the future.
-Choose one of the following answers: Strongly disagree, Disagree, Neither agree or disagree, Agree or Strongly Agree.
-D.5. Part 5 – Background
-Q1.
-How much do you agree with the following statement:
-  • I trust password managers (with my passwords).
-  • I think Password Managers are, overall, difficult to use.
-  • I feel my password are safer in a PM.
-  • I feel Password Managers are, overall annoying to use.
-  • I am willing to pay for software products.
-  • I understand how Password Managers work.
-Choose one of the following answers: Strongly disagree, Disagree, Neither agree or disagree, Agree or Strongly Agree.
-Q2.
-When using a browser do you allow the browser to remember your password?
-Q3.
-Do you know what formal verification is? Please explain if you do.
-Q4.
-How do you describe your gender identity?
-Choose one of the following answers: Male; Female; Non-binary; Prefer not to say; Other.
-Q5.
-How old are you?
-Choose one of the following answers: 18-24 years old; 25-34 years old; 35-44 years old; 45-54 years old; Above 55.
-Q6.
-What devices do you use daily?
-Choose one, or more, of the following answers: Laptop; Desktop; Smartphone; Table; Other.
-Q7.
-What is the highest degree you have completed? (If currently enrolled, highest degree received.)
-Choose one of the following answers: Less than a high school diploma High school degree or equivalent; Bachelor’s degree; Master’s degree; Doctorate degree.
-Appendix E Final Questionnaire
-The final questionnaire was used as the basis for a semi-structured interview where participants were encouraged to think-aloud about the questions.
-Most Likert Scale survey answers were accompanied by the participants’ reasons for their answers.
-E.1. Part 1 – SUS
-This section was composed of the standard SUS Questionnaire.
-E.2. Part 2 – Knowledge about PMs
-1.
-What does this icon symbolize (in the context of PROJECT_NAME)?
-See Figure 2.
-2.
-How much do you agree with the following statements:
-  • I feel my password are less safe in a PM.
-  • I feel I could trust in a Password Manager to save my passwords for me.
-  • I know how a Password Manager works.
-  • It is more convenient to use a Password Manager than to memorize passwords.
-  • I think Password Managers are, overall, difficult to use.
-  • If I were to use a Password Manager it would not be important for me to understand how the Password Manager works.
-Choose one of the following answers: Strongly disagree, Disagree, Neither agree or disagree, Agree or Strongly Agree.
-2.
-When using a PM, how important is:
-  • Ease of use.
-  • Support material (tutorials, help pages).
-  • Security and Privacy/Feeling that my data is secure.
-  • Having formally verified features.
-  • Price.
-  • Understanding how the Password Manager works.
-  • Being open source (i.e. the code is open for everyone to see).
-  • The company that made and manages the PM.
-Choose one of the following answers: Not at all important, Slightly Important, Important, Fairly Important or Very Important.
-E.3. Part 2 – Impact of Formal Verification
-1.
-How much do you agree with the following statements:
-  • I feel my password are less safe in a formally verified PM.
-  • I feel I could trust in a formally verified Password Manager to save my passwords for me.
-  • I know how a formally verified Password Manager works.
-  • I prefer Password Managers that are not formally verified.
-  • I would be willing to pay more for a software product if I knew it was formally verified.
-  • I would not trust a product just because it was formally verified.
-  • I do not value the use of Formal Verification in software products.
-Choose one of the following answers: Strongly disagree, Disagree, Neither agree or disagree, Agree or Strongly Agree.
-2.
-Do you know what formal verification is? Please explain.
-3.
-In your understanding, what is formal verification?
-4.
-What specific features were formally verified in the Password Manager you used? What is the used of formal verification in the features you mentioned.
-E.4. Part 3 – Perception of Safety and Final Questions
-1.
-After using PROJECT_NAME’s Password Manager how much do you agree with the following statements:
-  • I felt safe using the PM.
-  • I felt my password were secure in the PM.
-  • I would be more willing to use a Password Manager in the future if I knew it was formally verified.
-Choose one of the following answers: Strongly disagree, Disagree, Neither agree or disagree, Agree or Strongly Agree. Justify your answers.
-2.
-Will you use/continue to use a Password Manager in the future?
-Appendix F Scenarios descriptions
-Table 4. Scenarios descriptions.
-PM Feature Name of Scenario Scenario
-Password Generator Scenario 1 Unpredicta- bility Imagine that you use a Password Manager to generate several new random passwords for services that you use (e.g.
-Twitter, Facebook, internet banking).
-After generating a few passwords, you notice that all the generated passwords share a common pattern: they all start and end with the same set of letters.
-Because of this, the password was predicted by an attacker and now they have this password.
-Scenario 2 Policy Compliance Imagine that you are creating a new account on a website (e.g.
-Twitter, Facebook).
-To increase security, you ask the Password Manager to generate a password with 7 characters and with at least 2 numbers.
-However, the password generated does not include any numbers.
-Password Storage Scenario 3 Vault exposure Imagine that you are using a Password Manager to store your passwords and you discover that an attacker was able to access all your passwords saved in the Password Manager without knowing the primary password.
-Primary Password Scenario 4 Primary Password Exposure Imagine that you log in to your Password Manager with your primary password, but it is exposed and, as a result, an attacker learns your primary password.
-Autofill Scenario 5 Autofill The autofill feature of a Password Manager automatically fills your username and password in a website’s login page to make logging in easier.
-Imagine that you want to login in a website with a password that is saved in the Password Manager.
-The autofill feature fills in the password field and you log in.
-After this, you discover that an attacker was able to discover your password to that website because the autofill feature left the password exposed.
-Clipboard Scenario 6 Clipboard Clearing Imagine that you want to log in to a website.
-To log in you decide to copy the password from the Password Manager and paste it into the website to log in.
-When you copy the password it can be pasted anywhere.
-But, to prevent an attack, 30 seconds after you copy the password, the Password Manager should clear it so an attacker can’t paste it again and discover it.
-However, this doesn’t happen and, as a result, an attacker learns your password.
-Synchroniza- tion across devices Scenario 7 Synchroni- zation The Password Manager that you use allows you to keep a backup of all your passwords.
-Imagine now that an attacker deletes all your passwords, and when you try to restore them with the backup, you find that not all passwords were backed up and, as such, there are some passwords that you lost.
-Scenario 8 Ransomware/ Deleting your vault The Password Manager that you use allows you to keep a backup of all your passwords.
-Imagine now that you lose access to your computer and need to access your passwords in the Password Manager from another location.
-However, you discover that not all passwords were backed up and, as such, there are some passwords that you lost.
-Appendix G Second Study Survey
-The survey used in the second study is divided in four main parts: Password Managers (§G.2), Mathematically Correct Password Manager (§G.3), Scenarios (§G.4), and Background (§G.5).
-The survey starts with the participants going through the informed consent form (§G.1).
-G.1. Consent
-Thank you for your interest in participating in the survey ”Users’ willingness to use Password Managers”.
-Below, you find information about this research project, conditions for participation, and handling of the collected data.
-Please read everything carefully.
-If you agree and wish to participate in this study, please confirm your consent below.
-General information about the research project
-This study is conducted in the context of the PROJECT_NAME. The goal of this survey is to investigate password managers and factors that impact their use.
-In this study, you will be asked to complete a questionnaire about Password Managers.
-To this end, we will ask you to indicate how strongly you agree or disagree with statements and scenarios.
-We will ask you some demographic data.
-The survey will take approximately 10 minutes to complete.
-There are no specific prerequisites for participating in this study.
-Please note that you must be at least 18 years old to participate in this study.
-You can earn £1.65 for your participation.
-No particular burdens or damages from participating in this research project are to be expected.
-Voluntariness
-Your participation in this research project is voluntary. You can revoke your consent to participate at any time, without providing reasons, and without any disadvantages.
-Responsible management of the research project
-If you have any questions regarding the research project or if you wish to make use of your right of revocation, please contact:
-NAME and EMAIL ADDRESS
-There are 25 questions in this survey.
-G.2. Part 1 – Password Managers
-Please read the following scenario carefully.
-Imagine that you saw a news story about the use of Password Managers.
-In the news, they stated that Password Managers are programs that store and generate passwords.
-They stated that Password Managers can help increase your password security online.
-With a Password Manager, you save all your passwords in a secure vault and only have to remember one password — the primary password.
-Imagine now that you were thinking about using a Password Manager.
-How much do you agree with the following options:
-I would be more willing to use a Password Manager that…
-  • has support materials like tutorials or help pages
-  • is mathematically correct, that is, its features are as trustworthy as a mathematical proof
-  • is inexpensive
-  • is free
-  • is certified by PMSG (Password Manager Security Group)
-  • is made by a trustworthy company that is familiar to you
-  • please select the option “Strongly disagree”
-  • is easy to use for first-time and beginner users
-Please choose the appropriate response for each item: Strongly disagree, Disagree, Neither agree or disagree, Agree or Strongly agree.
-G.3. Part 2 - Mathematically Correct Password Manager
-Depending of the answer to the previous question there are three possible questions that could be asked in this part of the survey.
-G.3.1. Option A
-Only answer this question if answer was ‘Agree’ or ‘Strongly agree’ at question How much do you agree with the following options: I would be more willing to use a Password Manager that… (is mathematically correct, that is, its features are as trustworthy as a mathematical proof).
-In the previous question you stated you felt more willing to use a Password Manager that is mathematically correct, that is, its features are as trustworthy as a mathematical proof.
-Please state your reasons.
-G.3.2. Option B
-Only answer this question if the following conditions are met: Answer was ‘Strongly disagree’ or ‘Disagree’ at question How much do you agree with the following options: I would be more willing to use a Password Manager that… (is mathematically correct, that is, its features are as trustworthy as a mathematical proof).
-In the previous question, you stated you felt less willing to use a Password Manager that is mathematically correct, that is, its features are as trustworthy as a mathematical proof.
-Please state your reasons.
-G.3.3. Option C
-Only answer this question if the following conditions are met: Answer was ‘Neither agree or disagree’ at question How much do you agree with the following options: I would be more willing to use a Password Manager that… (is mathematically correct, that is, its features are as trustworthy as a mathematical proof).
-In the previous question, we asked you if you felt more willing to use a Password Manager that is mathematically correct, that is, its features are as trustworthy as a mathematical proof.
-You stated that you neither agree nor disagree with this statement.
-Please state your reasons.
-G.4. Part 3 – Scenarios
-Remember that Password Managers are programs that store and generate passwords.
-With a Password Manager, you save all your passwords in a secure vault and only have to remember one password — the primary password.
-In this section we’ll present you with scenarios and for each we ask how much do you agree with the following:
-I would stop using a Password Manager if this scenario happened.
-Please read each scenario carefully.
-G.4.1. Scenario A
-I would stop using a Password Manager if this scenario happened:
-Imagine that you use a Password Manager to generate several new random passwords for services that you use (e.g.
-Twitter, Facebook, internet banking).
-After generating a few passwords, you notice that all the generated passwords share a common pattern: they all start and end with the same set of letters.
-Because of this, the password was predicted by an attacker and now they have this password.
-Choose one of the following answers: Strongly disagree, Disagree, Neither agree or disagree, Agree or Strongly Agree.
-G.4.2. Scenario B
-I would stop using a Password Manager if this scenario happened:
-Imagine that you are creating a new account on a website (e.g.
-Twitter, Facebook).
-To increase security, you ask the Password Manager to generate a password with 7 characters and with at least 2 numbers.
-However, the password generated does not include any numbers.
-Choose one of the following answers: Strongly disagree, Disagree, Neither agree or disagree, Agree or Strongly Agree.
-G.4.3. Scenario C
-I would stop using a Password Manager if this scenario happened:
-Imagine that you are using a Password Manager to store your passwords and you discover that an attacker was able to access all your passwords saved in the Password Manager without knowing the primary password.
-Choose one of the following answers: Strongly disagree, Disagree, Neither agree or disagree, Agree or Strongly Agree.
-G.4.4. Scenario D
-I would stop using a Password Manager if this scenario happened:
-Imagine that you log in in your Password Manager with your primary password, but it is exposed and, as a result, an attacker learns your primary password.
-Choose one of the following answers: Strongly disagree, Disagree, Neither agree or disagree, Agree or Strongly Agree.
-G.4.5. Scenario E
-I would stop using a Password Manager if this scenario happened:
-The autofill feature of a Password Manager automatically fills your username and password in a website’s login page to make logging in easier.
-Imagine that you want to log in in a website with a password that is saved in the Password Manager.
-The autofill feature fills in the password field and you log in.
-After this you discover that an attacker was able to discover your password to that website because the autofill feature left the password exposed.
-Choose one of the following answers: Strongly disagree, Disagree, Neither agree or disagree, Agree or Strongly Agree.
-G.4.6. Scenario F
-I would stop using a Password Manager if this scenario happened:
-Imagine that you want to log in to a website.
-To log in you decide to copy the password from the Password Manager and paste it into the website to log in.
-When you copy the password it can be pasted anywhere.
-But, to prevent an attack, 30 seconds after you copy the password, the Password Manager should clear it so an attacker can’t paste it again and discover it.
-However, this doesn’t happen and, as a result, an attacker learns your password.
-Choose one of the following answers: Strongly disagree, Disagree, Neither agree or disagree, Agree or Strongly Agree.
-G.4.7. Scenario G
-I would stop using a Password Manager if this scenario happened:
-The Password Manager that you use allows you to keep a backup of all your passwords.
-Imagine now that an attacker deletes all your passwords, and when you try to restore them with the backup, you find that not all passwords were backed up and, as such, there are some passwords that you lost.
-Choose one of the following answers: Strongly disagree, Disagree, Neither agree or disagree, Agree or Strongly Agree.
-G.4.8. Scenario H
-I would stop using a Password Manager if this scenario happened:
-The Password Manager that you use allows you to keep a backup of all your passwords.
-Imagine now that you lose access to your computer and need to access your passwords in the Password Manager from another location.
-However, you discover that not all passwords were backed up and, as such, there are some passwords that you lost.
-Choose one of the following answers: Strongly disagree, Disagree, Neither agree or disagree, Agree or Strongly Agree.
-G.4.9. Scenario I
-I would stop using a Password Manager if this scenario happened:
-To get results that are relevant, all participants need to pay attention to the scenarios in the survey.
-This question is to make sure you are paying attention to the study, so please choose the answer ”strongly agree“.
-Choose one of the following answers: Strongly disagree, Disagree, Neither agree or disagree, Agree or Strongly Agree.
-G.5. Part 4 – Background
-1.
-Are you familiar with the concept of formal verification?
-Choose one of the following answers: Yes, No or I’m not sure
-2.
-If you are familiar with the concept of formal verification, in your understanding, what is formal verification?
-3.
-Did you know what a password manager was before this study?
-Choose one of the following answers: Yes, No or I’m not sure
-4.
-Have you used a password manager in the past? Choose one of the following answers:
-  • Yes, and I currently use one;
-  • Yes, but not anymore;
-  • No;
-  • I don’t know
-5.
-Which password manager did you use?
-Only answer this question if answer was ’Yes, and I currently use one’ or ’Yes, but not anymore’ at question Have you used a password manager in the past?.
-6.
-How do you remember your passwords? Select all that apply.
-Check all that apply:
-  • I save my passwords in a password manager
-  • I write my passwords down on paper
-  • I write my passwords in documents (e.g. text files or excel sheets)
-  • I remember my passwords from memory
-  • Other:
-7.
-When you use a login in a website sometimes a pop-up appears asking you to save the password for that website (this pop-up can appear, for example, in Chrome, Firefox, Safari, Apple Keychain).
-Do you save your passwords when prompted?
-Choose one of the following answers:
-  • Yes, I always save my passwords when asked;
-  • Yes, I sometimes save my passwords when asked;
-  • No, but I’ve saved passwords when asked in the past;
-  • No, I never save my passwords when asked;
-  • I don’t know
-G.6. Part 5 - Demographics
-1.
-How do you describe your gender identity?
-Choose one of the following answers: Male; Female; Non-binary; Prefer not to say; Other.
-2.
-How do you describe your race or ethnic identity?
-Check all that apply: Black; White; Hispanic; Asian; Native American; Prefer not to say; Other.
-Your survey responses have been recorded.
-Thank you for completing this survey.
-Appendix H Codebook used in the second study
-Table 5. Codebook used in second user study
-Code Description of code
-Frequency
-of Code
-Extra security
-The participants perceive the Password Manager more positively because it is more secure or with more safety. But they
-do not specify in what way it is more secure.
-47
-Extra security: password
-generation
-The participants perceive the Password Manager more positively because it is more secure. They specify that the password
-generation is more secure.
-26
-Extra security: protection
-from unwanted access
-The participants perceive the Password Manager more positively because it is more secure. They specify that it can prevent
-attacks from bad third parties OR specify that it is harder to hack the password.
-29
-Extra security: secure
-storage
-The participants perceive the Password Manager more positively because its more secure. They specify that the storage is
-more secure. For example there is low chance of password leaks
-18
-Use of mathematics The participant mentions that the Password Manager employ scientific or mathematically principles. 67
-Just seems better The participant believes that a Password Manager with formal verification ”just seems better” with no specific reason. 8
-Trustworthy
-The participant thinks they would trust this Password Manager more and that is why they would like to use a mathematical
-correct PM.
-39
-More assurances The behavior of the Password Manager is more certain. This may not have anything to do with security (e.g. less bugs). 14
-Distrust of Password Managers The participant does not trust in Password Managers in general. 5
-Not relevant The participant finds that even if a Password Manager is mathematically correct this is not relevant for them. 12
-Lack of understanding
-Lack of understanding about what mathematically correct means makes users not sure if it’s good OR about
-what is being proposed in the question and survey.
-30
-Not related to verification
-or mathematics.
-The participant response does not seem related to verification or mathematics. 5
-Answer not relevant The answer is not relevant to this study. 2
+Moreover, our work has shed light on a previously mostly unexplored area of research, combining formal verification with usable security.
+We argue that many paths for future work could be explored: virtually any domain where formal methods could be applied could benefit from studying user perception on the subject.
+{credits}
+7.0.1 Acknowledgements
+This work was financed by National Funds through the FCT - Fundação para a Ciência e a Tecnologia, I.P.
+(Portuguese Foundation for Science and Technology) under grant PRT/BD/153739/2021 and project UIDB/50021/2020 (DOI: 10.54499/UIDB/50021/2020).
+Work done in the context of the PassCert project, a CMU Portugal Exploratory Project funded by FCT with reference CMU/TIC/0006/2019.